### v1

# Beyond Semantics: Rediscovering Spatial Awareness in Vision-Language Models

Abstract
Vision-Language Models (VLMs) excel at identifying and describing objects but struggle with spatial reasoning such as accurately understanding the relative positions of objects. Inspired by the dual-pathway (ventral-dorsal) model of human vision, we investigate why VLMs fail spatial tasks despite strong object recognition capabilities. Our interpretability-driven analysis reveals a critical underlying cause: vision embeddings in VLMs are treated primarily as semantic “bag-of-tokens,” overshadowing subtle yet crucial positional cues due to their disproportionately large embedding norms. We validate this insight through extensive diagnostic experiments, demonstrating minimal performance impact when token orders or fine-grained spatial details are removed. Guided by these findings, we propose simple, interpretable interventions, including normalizing vision embedding norms and extracting mid-layer spatially rich features, to restore spatial awareness. Empirical results on both our synthetic data and standard benchmarks demonstrate improved spatial reasoning capabilities, highlighting the value of interpretability-informed design choices. Our study not only uncovers fundamental limitations in current VLM architectures but also provides actionable insights for enhancing structured perception of visual scenes. Website is at: user074.github.io/respatialaware/
Figure 1: Permutation Test: Original (top) vs. randomly permuted vision tokens (bottom). Despite losing spatial ordering, the LLM accurately responds to the prompt “Describe the image,” demonstrating strong robustness and a notable “bag-of-tokens” tendency. Token embeddings are visualized using cosine similarity relative to a reference token.
1 Introduction
Humans process vision through two distinct pathways: the ventral stream, which identifies objects, and the dorsal stream, which encodes spatial relationships [10]. The dorsal stream, located in the parietal lobe, is essential for processing spatial relationships, allowing us to understand where objects are and how they relate to each other. Interestingly, patients with damage to the parietal lobe exhibit asymmetric visual processing: they retain the ability to recognize and describe objects but lose spatial reasoning capabilities [10].
Recent observations suggest that Vision-Language Models (VLMs) exhibit a similar asymmetry. Despite remarkable performance in object-centric (ventral) tasks such as object recognition and image captioning[11, 2, 12], VLMs systematically struggle with even fundamental spatial reasoning queries (e.g., distinguishing “left” from “right”) [19, 33]. For example, when asked, “Is the apple to the left of the cup?”, current state-of-the-art VLMs can accurately identify both objects, yet frequently fail to correctly infer their spatial relationship. This leads to a critical interpretability question: Why do VLMs systematically fail at spatial reasoning even though they accurately recognize objects?
One possible reason is the inherent treatment of visual inputs in VLMs as a “bag-of-tokens”, analogous to the “bag-of-words” representation in NLP, where token order is disregarded [13, 38]. In this kind of representation, vision embeddings lose intrinsic spatial structure, significantly limiting relational awareness. However, the underlying interpretability mechanisms behind this phenomenon remain poorly understood.
We propose that adopting the ventral-dorsal analogy provides a useful interpretability heuristic to analyze how VLMs process what versus where information. Specifically, we hypothesize that spatial reasoning failures are exacerbated not merely by insufficient positional encoding but crucially by embedding norm suppression—the phenomenon where large embedding magnitudes overwhelm positional signals in attention mechanisms.
To rigorously validate these hypotheses, we conduct a comprehensive interpretability-driven analysis. Building upon these interpretability insights, we propose a minimal yet interpretable solution, namely, embedding norm regulation, which is aimed at explicitly restoring 2D spatial awareness in VLMs. Through controlled experiments, we demonstrate that embedding norm normalization effectively enhances spatial reasoning without compromising semantic understanding.
In summary, our contributions include:
  1. 1.
    An interpretability-driven theoretical explanation for VLMs’ spatial reasoning failures, rooted in embedding norm suppression.
  2. 2.
    Empirical interpretability evidence validating this theoretical insight through systematic permutation tests, embedding analysis, and attention visualization.
  3. 3.
    A simple, interpretable intervention that explicitly addresses spatial reasoning deficiencies, significantly improving performance on spatially demanding tasks.
The paper is organized as follows. After the introduction in Section 1, related work on VLMs, their intepretability and position encoding in Transformers is discussed in Section 2. Section 3 provides a diagnosis of the “Bag of Tokens” issue with token permutation and spatial compression. Detailed empirical and theoretical analyses are provided in Section 4 on spatial norm suppression. Section 5 proposes a controlled study in restoring spatial awareness of VLMs by designing a synthetic spatial benchmark. Experimental results on both synthetic and standard benchmarks are provided in Section 6. Section 7 analyzes how the combination of embedding normalization and multi-layer feature extraction influences the attention mechanisms of VLM. Section 8 discusses the findings as well as the limitations of the work, and Section 9 provides a few concluding remarks.
2 Related Work
Vision-Language Models Recent Vision-Language Models (VLMs) combine pretrained vision encoders and large language models (LLMs) to perform image understanding tasks, such as image captioning and visual question answering [23, 24, 20, 1, 41, 28, 7, 15]. Many approaches leverage pretrained visual encoders, notably CLIP [27], and subsequently introduce an adapter module to integrate vision embeddings into pretrained LLMs, such as Vicuna [6], enabling coherent textual outputs [23, 41]. While these methods achieve impressive performance in descriptive tasks, they often struggle with spatial reasoning [33, 38]. To address spatial limitations, recent efforts have focused on augmenting training datasets with extensive spatial annotations [32, 3, 5, 34]. In contrast, our approach does not rely on additional spatial labels or modifications to the original training data; instead, we introduce representation-level interventions, including embedding norm normalization and mid-layer feature extraction, which demonstrate significant improvements in VLMs’ spatial reasoning capabilities.
VLM Interpretability Interpretability of Vision-Language Models has attracted increasing attention, with the aim of explaining the internal mechanisms driving their successes or failures. Neo et al. [26] show that vision tokens in LLaVA [23] have strong semantic properties. Concurrently, multiple studies have highlighted the consistent failures of VLMs to comprehend simple spatial relations such as “above,” “below,” “left,” and “right” [38, 19, 33, 4]. These findings underscore a critical need for enhanced interpretability and robust spatial grounding in VLM architectures. Our work investigates the underlying reasons why spatial information is overshadowed in LLM, offering minimal yet effective interventions that notably strengthen the spatial reasoning ability.
Position Encoding in Transformers Positional encoding is essential in Transformer architectures to prevent degenerating into bag-of-words representations and to effectively model sequence orders [9, 35]. Initial efforts employed absolute positional encodings, later evolving into relative positional encodings [29] and, more recently, advanced mechanisms such as Rotary Positional Embeddings (RoPE) to better capture complex sequence dependencies [30]. Positional encoding has proven essential for modeling both language and visual data in Transformers [8]. Despite these developments, the specific impact of positional encodings within VLMs—where vision tokens interact with textual tokens in an LLM decoder—remains underexplored. Recent work [37] attributed hallucination issues to RoPE’s long-term decay property. Complementing these insights, our work empirically and theoretically illustrates how disproportionately large embedding norms of vision tokens can overshadow positional cues, hindering spatial reasoning even when advanced positional methods like RoPE are employed.
3 Diagnosing the Issue: Bag of Tokens
Dataset Original Permutation Difference
VQAv2 78.2 77.35 -0.85
POPE 87.3 87.10 -0.2
GQA 61.36 58.62 -2.74
CV-Bench 2D 56.59 56.26 -0.33
Table 1: Impact of random vision token permutation on spatial reasoning accuracy (%).
VLMs demonstrate excellent capabilities in recognizing objects within images (ventral tasks) yet consistently struggle with spatial reasoning (dorsal tasks). This disparity suggests that VLMs might neglect spatial cues in their internal representations. Open-source VLMs such as LLaVA [23] typically fuse vision embeddings with textual embeddings in transformer-based language models, relying critically on positional embeddings to encode token order. We hypothesize that in multimodal contexts, positional embeddings might become ineffective or suppressed, causing a loss of meaningful spatial structure. To investigate this, we perform two diagnostic interpretability experiments: (1) Token Permutation Test: If positional order significantly contributes to spatial reasoning, randomizing token order should substantially degrade model performance. We want to know how much the vision token order matters. (2) Spatial Compression Study: If local spatial information is crucial, drastically reducing the number of vision tokens should cause significant performance deterioration. We want to know how much VLM really uses local spatial information. Our experiments indicate that VLMs remain largely robust to these two perturbations, reinforcing the hypothesis of a “bag-of-tokens” behavior.
3.1 Token Permutation Test
In standard NLP transformers, positional embeddings are essential for maintaining token order. Similarly, if VLMs depend heavily on positional embeddings for spatial reasoning, randomizing vision token order should substantially impair spatial reasoning performance. We hypothesize that if token order carries meaningful spatial information, random shuffling should severely degrade performance.
Experiment Setup: We use LLaVA 1.5 7B from [23, 24] to conduct the test. We randomly permute vision token embeddings obtained from the vision encoder and projection layers before feeding them into the LLM, and the performance is then compared against the original unpermuted baseline.  Figure 1 illustrates the experiment pipeline. We evaluate performance on datasets including VQAv2[11], POPE [21], GQA [16], and CV-Bench[32].
Result: As shown in  Table 1, random token permutation leads only to minor performance drops (ranging from 0.2% to 2.74%). This surprisingly small decline reveals that VLMs exhibit significant order-invariance, suggesting a bag of tokens tendency of VLMs. This is summarized as the following “Intepretabality Insight”:
Interpretability Insight No. 1:
VLMs exhibit limited sensitivity to vision token order, reflecting a strong “bag-of-tokens” tendency.
Figure 2: Performance impact of vision token compression on standard benchmarks (GQA, CV-Bench 2D, and POPE). Only minor accuracy degradation occurs, even under extreme token compression (down to a single token).
3.2 Spatial Compression Study
Recent work showed VLMs focus on object semantic tokens[26]. To further probe the use of spatial information by VLMs, we progressively compress the vision tokens, effectively removing fine-grained spatial details including detailed positional and geometric relationships among image patches. If spatial details are crucial for evaluation, performance should sharply decline as spatial resolution decreases. Conversely, stable performance despite extreme compression may indicate either that the model relies predominantly on semantic features or that existing benchmarks do not effectively challenge spatial reasoning capabilities.
Experiment Setup: Using the LLaVA 1.5 7B model, we systematically reduce the number of vision tokens from the original 576 down to just 1. Models were retrained for token counts (256, 64, 16, and 1) using average pooling after the MLP projection layer. This method severely limits local spatial details while preserving some semantic content. Performance impacts are evaluated across various benchmarks.
Result: As illustrated in  Figure 2, reducing the vision tokens from 576 down to just 1 for standard benchmarks results in a minor accuracy drop (-8.5% the worst case). Despite losing approximately 99.8% of spatial resolution, the model’s performance remains robust. This outcome suggests two interrelated explanations: first, VLMs may primarily rely on semantic cues rather than local spatial details when solving typical benchmark tasks; second, the current benchmarks themselves might not effectively test the model’s capacity for detailed spatial reasoning. The limited sensitivity of these benchmarks to extreme spatial compression could reflect both model behavior and benchmark limitations. This leads us to the following insight:
Interpretability Insight No. 2:
For current benchmarks, VLMs can predominantly rely on “ventral” semantics rather than “dorsal” spatial cues to have a decent performance.
4 Why Does This Happen? Analysis of Embedding Norm Suppression
In earlier experiments, we observed that spatial information in VLMs tends to be minimal or overshadowed by strong global semantic signals. Here, we further investigate the underlying cause by analyzing embedding norms, specifically examining how large vision embedding magnitudes overshadow positional encodings within attention mechanisms.
4.1 Embedding Norm Analysis
Motivation Positional embeddings were initially introduced into Transformer architectures to maintain token order and avoid the “bag-of-words” effect [35]. While VLMs also utilize positional encodings during training, we hypothesize these encodings become ineffective if vision embeddings possess disproportionately large magnitudes, thus dominating positional signals during attention computations.
To investigate this hypothesis, we conduct an empirical embedding analysis using the COCO validation dataset (5000 image-text pairs) [22]. Specifically, we extract vision and text embeddings after MLP projection but before input into the language model and compute the L2 norms for each token type. Figure 3 illustrates the embedding norm distributions.
Figure 3: Distribution of L2 norms for vision and text tokens in COCO validation dataset (log scale). Vision token norms range between $10^{1}$ and $10^{3}$, while text token norms range between $3\times 10^{-1}$ and $10^{0}$.
As shown in Figure 3, vision embeddings consistently have 1-2 orders of magnitudes greater than text embeddings but can reach up to 3 orders of magnitude greater at its upper range. Although directional alignment (e.g., cosine similarity) typically receives primary focus in multimodal alignment due to inner product [14, 17, 25], such large magnitude discrepancies may suppress positional signals significantly.
4.2 Theoretical Explanation
We provide a concise theoretical justification of how differences in embedding norms significantly impact the effectiveness of positional encodings in Transformer attention mechanisms. A detailed mathematical derivation is presented in the Appendix A but here is a summary.
Summary of Proof (detailed derivation in Appendix A). Consider a Transformer self-attention block employing Rotary Positional Embeddings (RoPE) [31]. RoPE applies orthonormal rotations to query and key vectors with dimension $d$ ($\mathbf{q}_{i},\mathbf{k}_{i}\in\mathbb{R}^{d}$), preserving their norms. After RoPE transformation, we denote these vectors as $\mathbf{q}_{i}^{\prime},\mathbf{k}_{i}^{\prime}\in\mathbb{R}^{d}$. The attention logits between token pairs are computed via the scaled dot-product:
$\mathrm{logits}_{ij}=\frac{\mathbf{q}_{i}^{\prime}\cdot\mathbf{k}_{j}^{\prime}}{\sqrt{d}}=\frac{\|\mathbf{q}_{i}\|\|\mathbf{k}_{j}\|\cos{\phi}}{\sqrt{d}}$ (1)
where $\phi$ denotes the positional angle difference. From this expression, we observe directly that larger embedding norms increase the scale of logits. When vision embedding norms substantially exceed text embedding norms (often by an order of magnitude), vision-to-text logits dominate softmax computations, resulting in disproportionately high attention scores. To maintain coherent textual outputs aligned with the language model’s training, the model implicitly learns to downscale these vision-related logits. This downscaling leads to empirically observed low attention weights for vision tokens, consistent with findings in recent studies [4]. Consequently, the positional sensitivity explicitly encoded by RoPE diminishes substantially, as lower attention weights inherently reduce the positional gradient (further detailed in Appendix A).
Interpretability Insight No. 3 The practical effectiveness of positional encoding mechanisms’ (such as RoPE) is severely limited by embedding magnitude discrepancies between modalities.
Normalizing or rescaling embedding norms thus emerges as a crucial intervention, restoring positional sensitivity and enabling positional embeddings to meaningfully influence attention computations. This insight extends broadly to multi-modal alignment scenarios: effective alignment between pretrained models should not rely solely on embedding similarities (such as cosine similarity) but must also carefully consider embedding norm differences. Otherwise, positional information can become suppressed, leading to minimal influence on model predictions.
5 Restoring Spatial Awareness: A Controlled Study
Our previous analyses indicated that spatial cues in VLMs are often overshadowed by strong semantic signals and disproportionately large embedding norms. To validate these findings and explore methods for enhancing spatial representations, we designed a controlled study employing two interpretability-informed strategies:
  1. 1.
    Synthetic Dataset: A carefully constructed two-dimensional (2D) spatial benchmark designed explicitly to eliminate semantic shortcuts, ensuring that spatial relationships alone drive task success.
  2. 2.
    Interpretability-Informed Model Adjustments: Targeted architectural modifications aimed at balancing positional cues and mitigating the overwhelming dominance of semantic embeddings.
Combining these controlled experimental conditions with interpretability-driven adjustments allows us to investigate the feasibility of reviving effective spatial reasoning—analogous to restoring a functional “dorsal stream”—in VLMs.
5.1 A 2D Synthetic Spatial Benchmark (2DS)
Standard image and text benchmarks often unintentionally provide semantic or contextual shortcuts, enabling models to leverage object co-occurrence or linguistic biases without genuinely understanding spatial relationships. We particularly hope to respond to the potential inadequacies in current benchmarks from our spatial compression study (Section 2). To isolate and rigorously evaluate spatial reasoning, we limit confounding factors such as real-world object frequencies and linguistic priors. Our benchmark specifically targets 2D spatial relationships directly influenced by positional embeddings.
We constructed a synthetic dataset that: (1) Ensures spatial cues are essential for correctly answering queries. (2) Isolates and systematically controls semantic aspects, enabling analysis of their contributions to spatial understanding. The dataset combines semantic attributes (color, shape, or both) with spatial properties (relative or absolute positions). We denote the 2D Synthetic Dataset as 2DS
Example Queries including:
  • “What color is at the bottom of the image?” (Color, absolute position)
  • “Is the circle below the square?” (Shape, relative position)
These queries illustrate that object attributes (color, shape) serve primarily as identifiers, with positional relationships being crucial for correct responses.
5.1.1 Dataset Construction Details
We created meta-categories containing images with 2, 3, 4, 5, and 6 objects. Within each category, we employed the same set of objects while randomly varying their spatial positions. Each meta-category comprises 100 images, resulting in a total of 500 images. For every image, we systematically asked questions spanning combinations of semantic (color, shape, color and shape) and spatial (absolute, relative) properties, yielding 3000 total questions. Note the emphasis is on spatial relationship reasoning.
Figure 4 provides representative examples, emphasizing the simplicity of semantic attributes and the centrality of spatial relationships.
Figure 4: Illustrative scenes from our 2DS Dataset. The left two images show two object arrangements, while the right two images illustrate more complex three and four object arrangements.
5.2 Interpretability-Informed Model Adjustments
Our interpretability analyses identified embedding norm suppression and insufficient usage of local geometric information as critical limitations to effective spatial reasoning in VLMs. Guided by these insights, we propose two straightforward and interpretable adjustments: (1) Vision Token Normalization and (2) Utilizing Intermediate-Layer Features.
Vision Token Normalization Our prior analyses revealed that vision embedding magnitudes significantly exceed those of positional embeddings, effectively suppressing positional signals during attention computations. To address this imbalance:
  • Norm Matching: We measured the distribution of typical text embedding norms (mean $\approx 0.83$, max $\approx 1.22$).
  • RMS Normalization: We applied RMS normalization [39] that is used in LLM [6] to vision encoder outputs to match them closely with textual embedding norms.
This normalization reduces the dominance of semantic signals, allowing positional cues to emerge more clearly.
Intermediate-Layer Features (Local Geometry) Current VLM architectures typically use highly aggregated, semantically-rich final-layer vision features from CLIP, which excel in semantic alignment but may lack fine-grained spatial details [36]. Recent studies suggest that intermediate layers of CLIP encoders retain richer local spatial information beneficial for VLMs [18]. Thus, we investigate whether directly incorporating earlier spatial features improves spatial reasoning.
Specifically, we extracted features from intermediate vision encoder layers (12, 16, 20, and 24), inherently containing more detailed spatial information reminiscent of the dorsal-stream processing pathway. These mid-layer features explicitly encode fine-grained local spatial information, potentially crucial for accurately answering purely spatial queries in our synthetic dataset.
However, an interpretability-informed ablation study (Section 7.2) reveals a semantic-spatial trade-off: while intermediate features significantly boost purely spatial task performance, they might also reduce the LLM’s incentive to develop explicit positional sensitivity within its attention mechanism.
6 Experiments
Figure 5: Accuracy comparison across varying numbers of objects. Our interpretability-informed adjustments yield consistent improvements, especially as spatial complexity increases.
We now evaluate the effectiveness of our interpretability-informed interventions, namely vision embedding normalization and intermediate-layer feature extraction, in enhancing spatial reasoning in VLMs. Our experiments cover results on both a controlled synthetic spatial benchmark and several standard vision-language benchmarks.
6.1 Experimental Setup
We use LLaVA 1.5 (7B) [23] as our baseline model, which is chosen for its open-source availability and transparent training details. We replicate LLaVA 1.5’s exact two-stage training process and dataset usage for fair comparison. Based on the interpretability insights discussed earlier, we evaluate two proposed variants:
  • + Normalize: Vision embeddings normalized according to the procedure detailed in Section 5.2.
  • + Normalize + Multilayer: Normalized vision embeddings augmented with intermediate-layer features to retain fine-grained spatial information ( Section 5.2).
6.2 Results on Synthetic Dataset
Table 2 summarizes accuracy across distinct spatial query categories, and Figure 6 illustrates performance across varying numbers of objects, clearly highlighting the effectiveness of our interventions.
Figure 6: Accuracy comparison across different categories of 2DS dataset. We see consistent improvements across categories.
Category LLaVA 1.5 + Normalize + Normalize + Multilayer
Color_abs. $\uparrow$ 79.60 88.00 (+8.40) 87.80 (+8.20)
Color_rel. $\uparrow$ 37.00 42.40 (+5.40) 47.20 (+10.20)
Shape_abs. $\uparrow$ 78.60 79.00 (+0.40) 82.20 (+3.60)
Shape_rel. $\uparrow$ 34.40 32.80 (-1.60) 40.80 (+6.40)
Shape_color_abs. $\uparrow$ 70.80 71.80 (+1.00) 80.20 (+9.40)
Shape_color_rel. $\uparrow$ 39.40 41.80 (+2.40) 50.60 (+11.20)
Overall Acc. $\uparrow$ 56.63 59.30 (+2.67) 64.80 (+8.17)
Table 2: Spatial reasoning accuracy (%) across 2DS categories. $\uparrow$ indicates higher is better. Values in parentheses show difference from LLaVA 1.5 baseline.
Figure 7: Visualization of self-attention patterns. We overlay the attention map on top of the image. The question for the model is under each row. We use the first target word of the response for attention map, for example the attention map is based on ’square’ and ’circle’ for top left rows. Entropy values are on top of each image. Red answers are wrong. Green answers are correct. Each image has entropy on top: lower means more concentrated higher means more diffused
Vision Embedding Normalization: Normalization alone improves overall accuracy from 56.63% to 59.30%, strongly supporting our interpretability insight that balancing embedding norms significantly enhances spatial information utilization. Notably, improvements are particularly pronounced in color-related spatial relationship queries.
Normalization + Intermediate-layer Features: This approach achieves the highest overall accuracy (64.80%, +8.17% improvement over baseline), underscoring the importance of explicitly incorporating local spatial details. This method substantially enhances accuracy, especially in shape and color combined relative position queries.
Both methods consistently outperform the original LLaVA 1.5 model, particularly as spatial complexity (the number of objects) increases (Figure 5), validating our hypothesis about the critical roles of embedding normalization and local spatial detail preservation.
6.3 Results on Standard Benchmarks
Category LLaVA 1.5 + Normalize + Normalize + Multilayer
VQAv2 $\uparrow$ 78.20 78.76 (+0.56) 79.17 (+0.97)
POPE $\uparrow$ 87.30 87.30 (0.00) 87.70 (+0.40)
GQA $\uparrow$ 61.46 62.04 (+0.58) 62.52 (+1.05)
CV-Bench2D $\uparrow$ 56.59 59.91 (+3.32) 58.69 (+2.10)
Table 3: Performance (accuracy %) on standard vision-language benchmarks. $\uparrow$ indicates higher is better. Values in parentheses show difference from LLaVA 1.5 baseline.
To assess generalization, we evaluate our methods on established benchmarks (Table 3) that are used in recent works [23, 32, 41, 24].
Both interpretability-informed interventions yield moderate yet consistent improvements. Normalization and Multilayer Normalization consistently achieves small but meaningful gains (up to 3.32%), confirming that improved spatial representation benefits overall reasoning without sacrificing semantic capabilities. Notably, normalization shows more pronounced gains in the CV-Bench 2D dataset, particularly for spatial relationship tasks.
7 Analysis
To deepen our interpretability-driven understanding, we analyzed how the combination of embedding normalization and earlier-layer feature extraction influences the attention mechanisms of VLMs.
7.1 Attention Visualization
We visualized the self-attention patterns from LLM to interpret the effects of each modification on spatial reasoning. Figure 7 provides representative attention maps across four illustrative queries. Each query compares three scenarios: the baseline (Original LLaVA 1.5), normalized embeddings, and normalized embeddings combined with multilayer features. We visualized the first text token of the target object in the answer, which are bold and italic words in Figure 7. More examples are in Appendix C.
Baseline (Original LLaVA 1.5): The baseline model shows diffuse and scattered attention patterns, possibly influenced by large-magnitude vision embeddings that overshadow positional cues. This diffuseness makes it difficult for the model to clearly focus on relevant spatial tokens, leading to suboptimal spatial reasoning.
Embedding Normalization: Embedding normalization leads to distinct and focused attention distributions, significantly increasing attention on relevant spatial tokens. Interestingly, this variant exhibits broader coverage of the attention map, suggesting that the model actively utilizes positional embeddings to locate informative spatial regions. The improved spatial focus directly contributes to enhanced performance in spatial reasoning tasks.
Embedding Normalization + Multilayer Features: Interestingly, the multilayer feature variant produces sparser, more selective attention distributions with lower attention magnitudes, similar visually to the original model but fundamentally different in effectiveness. Despite lower peak attention, this variant consistently performs better across most spatial tasks. We hypothesize that intermediate-layer vision embeddings inherently carry sufficient local spatial information, thus significantly reducing the need for explicit positional encoding at the LLM stage. Consequently, the model selectively focuses attention on regions already identified as informative by intermediate-layer embeddings without the need to check other regions.
7.2 Interpreting Attention Patterns and Performance
To quantitatively interpret these observations, we computed the normalized attention entropy (Table 4). We average all text tokens’ attention over the vision tokens to compute the entropy. Lower entropy indicates more targeted attention distributions, and high scores mean more diffused attention distributions. More examples are in Appendix C.
Vision Token Attention Entropy LLaVA 1.5 + Normalize + Normalize + Multilayer
2DS $\downarrow$ 0.76 0.90 0.72
COCO Val $\downarrow$ 0.81 0.89 0.72
CV-Bench2D $\downarrow$ 0.80 0.90 0.72
Table 4: Average attention entropy over different datasets. Lower entropy indicates more selective attention.
The normalized embeddings exhibit higher entropy compared to the baseline, reflecting a broader and more uniform exploration of spatial regions, yet with a clear focus on the target object. This suggests that embedding normalization encourages the LLM to explicitly explore multiple spatial positions to extract positional information. In contrast, the multilayer normalized variant demonstrates the lowest entropy, underscoring the model’s confidence in selectively attending to a few precise spatial regions. This observation supports our hypothesis that intermediate-layer embeddings inherently encode robust spatial details, enabling the model to rely less on explicit attention-based positional signals to look at vision tokens. In addition, it suggests that the improvements of the Normalize and Multilayer Normalize are two different mechanisms: the former is based on LLM attention, and the later is based on vision encoder spatial information.
Interpretability Insight No.4.
This analysis underscores a fundamental interpretability trade-off: richer spatial information embedded earlier in the model pipeline reduces the dependence on explicit positional encoding effect at later stages. Consequently, LLM attention becomes sparser yet highly targeted, reflecting enhanced target in spatial information rather than a failure to attend effectively.
Providing early-layer features from vision encoders potentially offers a shortcut for LLMs to directly leverage spatial cues without extensive positional encoding training. Future research should explore strategies to optimally balance intermediate-layer spatial detail and positional encodings.
8 Discussion
Our experiments reveal that spatial reasoning in VLMs is often overshadowed by strong semantic cues and disproportionately large vision token norms. Normalizing embedding magnitudes and incorporating intermediate-layer vision features partially restore a dorsal-stream-like spatial awareness. The most pronounced improvements occur in our synthetic dataset, explicitly designed to emphasize spatial over semantic shortcuts.
This analysis represents an initial step toward deeper interpretability in VLMs. It highlights that not only dataset design but also detailed model analysis is crucial. Careful alignment of multi-modal representations, especially when leveraging pretrained models, is essential. Our findings provide actionable insights for achieving more effective modality alignment in future research.
Limitations Despite the promising results, our approach has several limitations: First, the proposed model adjustments are intentionally simple to illustrate our interpretability insights. The current normalization method is a brute-force matching to text embedding norms, and it can be overly simplistic. Second, there exists a fundamental trade-off regarding where spatial information is encoded. Normalization pushes LLMs to utilize positional signals explicitly, while multilayer normalization leverages inherent spatial details from vision encoders Finally, current evaluations focus primarily on 2D spatial reasoning tasks. In 3D, the larger magnitude might be more beneficial.
Future Directions. Future work should expand spatial benchmarks to robustly evaluate complex 3D reasoning and explore refined normalization methods that preserve semantic information while still promoting spatial awareness. Additionally, architectural modifications explicitly designed for balanced semantic and spatial reasoning could further advance VLM interpretability and effectiveness.
9 Conclusion
Through systematic diagnosis and interpretability-driven analysis of VLMs, we propose minimal yet impactful interventions that restore dorsal-stream-like spatial reasoning capabilities. Our insights highlight the critical role of embedding norms and layer-specific features, providing a clear interpretability pathway for future model improvements. As research increasingly seeks models capable of accurately understanding not only what objects exist but precisely where they are, our contributions lay a foundation for more spatially intelligent, interpretable, and effective VLMs.
Acknowledgments
The work is supported by the National Science Foundation (NSF) through Awards #2131186 (CISE-MSI), #1827505 (PFI), and the US Air Force Office of Scientific Research (AFOSR) via Award #FA9550-21-1-0082. The work is also supported by a College-wide Research Vision (CRV) Fund from the CCNY Provost’s Office.
Supplementary Material
Appendix A Theoretical Analysis of Large-Norm Effects on RoPE
Problem Setup
Consider a Transformer attention mechanism [35] employing Rotary Position Embeddings (RoPE)[30]. Each token $i$ generates a query $\mathbf{q}_{i}\in\mathbb{R}^{d}$ and a key $\mathbf{k}_{i}\in\mathbb{R}^{d}$. RoPE applies a position-dependent rotation $R(\cdot)$, transforming these vectors as follows:
$\mathbf{q}_{i}^{\prime}=R(\mathbf{q}_{i}),\quad\mathbf{k}_{i}^{\prime}=R(\mathbf{k}_{i}).$
For simplicity, assume $R(\cdot)$ preserves norms (this is standard for RoPE since it only rotates but does not change norm):
$\|\mathbf{q}_{i}^{\prime}\|=\|\mathbf{q}_{i}\|,\quad\|\mathbf{k}_{i}^{\prime}\|=\|\mathbf{k}_{i}\|.$
The attention logits are computed using scaled dot-product attention:
$\mathrm{logit}_{\text{txt},\text{vis}}=\frac{\mathbf{q}_{i}^{\prime}\cdot\mathbf{k}_{j}^{\prime}}{\sqrt{d}}$
1. Effects of Large Embedding Norms on Attention Weights
When vision embedding input into LLM directly without model parameter adjustments, we can generally assume the Query and Key of the vision embedding will scale accordingly. Consider two sets of embeddings: vision embeddings $\mathbf{q}_{\text{vis}}$, $\mathbf{k}_{\text{vis}}$ and text embeddings $\mathbf{q}_{\text{txt}}$, $\mathbf{k}_{\text{txt}}$. Empirically, we observed (as illustrated in the Section  4.1) that vision embeddings typically have substantially larger norms than text embeddings:
$\|\mathbf{q}_{\text{vis}}\|\approx M\|\mathbf{q}_{\text{txt}}\|,\mathbf{k}_{\text{vis}}\|\approx M\|\mathbf{k}_{\text{txt}}\|,\quad M\gg 1.$
Empirically, from our study $M$ can reach to the level of $10^{1}$ to $10^{2}$ range. Then based on Cauchy-Schwarz inequality we can simply drive following approximations for the attention logits for vision-to-text interactions:
$\mathrm{logit}_{\text{txt},\text{vis}}=\frac{\mathbf{q}_{\text{txt}}^{\prime}\cdot\mathbf{k}_{\text{vis}}^{\prime}}{\sqrt{d}}\leq\frac{\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{vis}}^{\prime}\|}{\sqrt{d}}\approx\frac{M\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}$
Similarly, for text-to-text interactions:
$\mathrm{logit}_{\text{txt},\text{txt}}\approx\frac{\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}$
Given the scale difference ($M\gg 1$), typically we can see that:
$\mathrm{logit}_{\text{txt},\text{vis}}\approx\frac{M\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}\gg\frac{\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}\approx\mathrm{logit}_{\text{txt},\text{txt}}$
It would follow that the softmax attention weights from vision tokens towards text tokens become disproportionately large if not adjusted:
$\alpha_{\text{txt},\text{vis}}$ $=softmax(\frac{\mathbf{q}_{\text{txt}}\cdot\mathbf{k}_{\text{vis}}}{\sqrt{d}})=\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{j}\exp(\mathrm{logit}_{\text{txt},j})}$ (2)
$=\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{j_{vis}}\exp(\mathrm{logit}_{\text{txt},j_{vis}})+\sum_{j_{txt}}\exp(\mathrm{logit}_{\text{txt},j_{txt}})}$ (3)
It is easy to see that the terms are dominated by the large $M$ discrepancies between the vision and text embeddings. However, practical model training implicitly enforces coherence conditioned primarily on textual tokens, forcing the model to scale down vision-query logits significantly. This downscaling results in empirically observed low attention weights for vision tokens. [4] observed the attention score of vision tokens are way much smaller compared to the attention weights of text tokens based on the score of the text token, and the score of vision token for most of layers are in $[0,0.2]$ range while text tokens are in $[0.8,1]$ range. It is reasonable to believe the model learned to scale down $\mathrm{logit}_{\text{txt},\text{vis}}$ significantly to balance out the large norm.
2. Impact on RoPE and Positional Encoding
The attention weight for token $j$ given query token $i$ is defined by the softmax function:
$\alpha_{\text{txt},\text{vis}}=\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{k}\exp(\mathrm{logit}_{\text{txt},k})}$
We analyze how sensitively the attention weight $\alpha_{\text{txt},\text{vis}}$ varies with respect to the positional angle $\phi$, introduced by the Rotary Position Embedding (RoPE). This sensitivity is given by the derivative:
$\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\frac{\partial}{\partial\phi}\left(\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{k}\exp(\mathrm{logit}_{\text{txt},k})}\right)$
Expanding this derivative explicitly with the quotient rule, we can denote $\exp(\mathrm{logit}_{\text{txt},\text{vis}})$ as $u$ and $\sum_{k}\exp(\mathrm{logit}_{\text{txt},k})$ as $v$, then we have:
$\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\frac{vu^{\prime}-uv^{\prime}}{v^{2}}$
then simplifying the exponential derivatives and factoring out $\alpha_{\text{txt},\text{vis}}$, we simplify it to:
$\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\alpha_{\text{txt},\text{vis}}\left(\frac{\partial\mathrm{logit}_{\text{txt},\text{vis}}}{\partial\phi}-\sum_{k}\alpha_{\text{txt},k}\frac{\partial\mathrm{logit}_{\text{txt},k}}{\partial\phi}\right)$
Key Insight
From this expression, we observe that the positional sensitivity (derivative) of the attention weight explicitly depends on the attention weight itself ($\alpha_{\text{txt},\text{vis}}$). Therefore, when the attention weight $\alpha_{\text{txt},\text{vis}}$ is small (as in the case of vision tokens after significant normalization), the derivative with respect to positional angle $\phi$ is correspondingly reduced, diminishing positional distinctiveness.
Mathematical Justification: The derivative of attention weights (softmax) w.r.t. positional angle $\phi$ is directly proportional to the attention weight itself: $\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\alpha_{\text{txt},\text{vis}}\left(\frac{\partial\mathrm{logit}_{\text{txt},\text{vis}}}{\partial\phi}-\sum_{k}\alpha_{\text{txt},k}\frac{\partial\mathrm{logit}_{\text{txt},k}}{\partial\phi}\right)$ Thus, lower attention weights inherently reduce positional sensitivity, explaining why downscaled vision embeddings exhibit diminished positional distinctiveness after normalization.
Appendix B Training Details
We follow the training recipe described in LLaVA 1.5 7B [23], consisting of two stages: pretraining and instruction tuning. All experiments utilize 8 Nvidia A100 GPUs. Learning rates for each training stage are detailed in Table 5, with all other hyperparameters and datasets identical to [23].
Model Pre-train Finetune
LLaVA 1.5 7B 1e-3 2e-5
+Normalize 1e-3 2e-5
+Normalize+Multilayer 1e-3 2e-5
Table 5: Learning rate of training
Appendix C Attention Entropy
We adapt the visualization method from [40] to analyze the LLM attention specifically on vision tokens. Attention maps correspond to particular, meaningful text tokens, as labeled above each image in Figure 8. This is a more refined version compared to average text tokens. Notably, we observe consistently lower entropy in the multilayer-normalized models.
Figure 8: More visualizations

### v2

@@
 Abstract
-Vision-Language Models (VLMs) excel at identifying and describing objects but struggle with spatial reasoning such as accurately understanding the relative positions of objects.
-Inspired by the dual-pathway (ventral-dorsal) model of human vision, we investigate why VLMs fail spatial tasks despite strong object recognition capabilities.
-Our interpretability-driven analysis reveals a critical underlying cause: vision embeddings in VLMs are treated primarily as semantic “bag-of-tokens,” overshadowing subtle yet crucial positional cues due to their disproportionately large embedding norms.
-We validate this insight through extensive diagnostic experiments, demonstrating minimal performance impact when token orders or fine-grained spatial details are removed.
-Guided by these findings, we propose simple, interpretable interventions, including normalizing vision embedding norms and extracting mid-layer spatially rich features, to restore spatial awareness.
-Empirical results on both our synthetic data and standard benchmarks demonstrate improved spatial reasoning capabilities, highlighting the value of interpretability-informed design choices.
-Our study not only uncovers fundamental limitations in current VLM architectures but also provides actionable insights for enhancing structured perception of visual scenes.
-Website is at: user074.github.io/respatialaware/
+Vision–Language Models (VLMs) excel at identifying and describing objects but often fail at spatial reasoning.
+We study why VLMs, such as LLaVA, under-utilize spatial cues despite having positional encodings and spatially rich vision-encoder features.
+Our analysis reveals a key imbalance: vision token embeddings have much larger norms than text tokens, suppressing LLM’s position embedding.
+To expose this mechanism, we developed three interpretability tools: (1) the Position Sensitivity Index, which quantifies reliance on token order, (2) the Cross-Modality Balance, which reveals attention head allocation patterns, and (3) a RoPE Sensitivity probe, which measures dependence on rotary positional embeddings.
+These tools uncover that vision tokens and system prompts dominate attention.
+We validated our mechanistic understanding through targeted interventions that predictably restore positional sensitivity.
+These findings reveal previously unknown failure modes in multimodal attention and demonstrate how interpretability analysis can guide principled improvements.
+Website is at: user074.github.io/respatialaware/.
 Figure 1: Permutation Test: Original (top) vs.
 randomly permuted vision tokens (bottom).
-Despite losing spatial ordering, the LLM accurately responds to the prompt “Describe the image,” demonstrating strong robustness and a notable “bag-of-tokens” tendency.
+Despite losing spatial ordering, the LLM accurately responds to the prompt “Describe the image,” demonstrating strong robustness and a notable “bag-of-tokens” tendency, losing spatial relationships.
 Token embeddings are visualized using cosine similarity relative to a reference token.
 1 Introduction
-Humans process vision through two distinct pathways: the ventral stream, which identifies objects, and the dorsal stream, which encodes spatial relationships [10].
+Humans process vision through two distinct pathways: the ventral stream, which identifies objects, and the dorsal stream, which encodes spatial relationships.
 The dorsal stream, located in the parietal lobe, is essential for processing spatial relationships, allowing us to understand where objects are and how they relate to each other.
-Interestingly, patients with damage to the parietal lobe exhibit asymmetric visual processing: they retain the ability to recognize and describe objects but lose spatial reasoning capabilities [10].
-Recent observations suggest that Vision-Language Models (VLMs) exhibit a similar asymmetry.
-Despite remarkable performance in object-centric (ventral) tasks such as object recognition and image captioning[11, 2, 12], VLMs systematically struggle with even fundamental spatial reasoning queries (e.g., distinguishing “left” from “right”) [19, 33].
-For example, when asked, “Is the apple to the left of the cup?”, current state-of-the-art VLMs can accurately identify both objects, yet frequently fail to correctly infer their spatial relationship.
-This leads to a critical interpretability question: Why do VLMs systematically fail at spatial reasoning even though they accurately recognize objects?
-One possible reason is the inherent treatment of visual inputs in VLMs as a “bag-of-tokens”, analogous to the “bag-of-words” representation in NLP, where token order is disregarded [13, 38].
-In this kind of representation, vision embeddings lose intrinsic spatial structure, significantly limiting relational awareness.
-However, the underlying interpretability mechanisms behind this phenomenon remain poorly understood.
-We propose that adopting the ventral-dorsal analogy provides a useful interpretability heuristic to analyze how VLMs process what versus where information.
-Specifically, we hypothesize that spatial reasoning failures are exacerbated not merely by insufficient positional encoding but crucially by embedding norm suppression—the phenomenon where large embedding magnitudes overwhelm positional signals in attention mechanisms.
-To rigorously validate these hypotheses, we conduct a comprehensive interpretability-driven analysis.
-Building upon these interpretability insights, we propose a minimal yet interpretable solution, namely, embedding norm regulation, which is aimed at explicitly restoring 2D spatial awareness in VLMs.
-Through controlled experiments, we demonstrate that embedding norm normalization effectively enhances spatial reasoning without compromising semantic understanding.
-In summary, our contributions include:
-  1. 1.
-    An interpretability-driven theoretical explanation for VLMs’ spatial reasoning failures, rooted in embedding norm suppression.
-  2. 2.
-    Empirical interpretability evidence validating this theoretical insight through systematic permutation tests, embedding analysis, and attention visualization.
-  3. 3.
-    A simple, interpretable intervention that explicitly addresses spatial reasoning deficiencies, significantly improving performance on spatially demanding tasks.
-The paper is organized as follows.
-After the introduction in Section 1, related work on VLMs, their intepretability and position encoding in Transformers is discussed in Section 2.
-Section 3 provides a diagnosis of the “Bag of Tokens” issue with token permutation and spatial compression.
-Detailed empirical and theoretical analyses are provided in Section 4 on spatial norm suppression.
-Section 5 proposes a controlled study in restoring spatial awareness of VLMs by designing a synthetic spatial benchmark.
-Experimental results on both synthetic and standard benchmarks are provided in Section 6.
-Section 7 analyzes how the combination of embedding normalization and multi-layer feature extraction influences the attention mechanisms of VLM.
-Section 8 discusses the findings as well as the limitations of the work, and Section 9 provides a few concluding remarks.
+Not surprisingly, patients with damage to the parietal lobe exhibit asymmetric visual processing: they retain the ability to recognize and describe objects but lose spatial reasoning capabilities (Goodale & Milner, 1992).
+Recent observations suggest that generative Vision-Language Models (VLMs) exhibit a similar asymmetry.
+Despite remarkable performance in object-centric (ventral) tasks such as object recognition and image captioning (Goyal et al., 2017; Antol et al., 2015; Gurari et al., 2018), VLMs struggle with even fundamental spatial reasoning queries (e.g., distinguishing “left” from “right”) (Kamath et al., 2023; Tong et al., 2024b).
+This leads to a critical interpretability question: Where does “space” live in VLMs, and why is it under-used?
+In this paper, we only focus on the common architecture that couples a pre-trained vision encoder with a pre-trained language model (LLM).
+In principle, both sides provide spatial signals: vision encoders carry layout through their feature maps and token order, while the LLM’s positional mechanism (RoPE) can make the decoder sensitive to the token order.
+Yet the combined system often behaves like a “bag of words”, mostly losing spatial relationships (Yuksekgonul et al., 2022).
+So the question is: Why does the integration fail?
+We hypothesize two causes: First, the RoPE is under-utilized.
+Second, the last few layers of the vision encoder might not contain enough spatial information.
+We first diagnose and show empirical evidence of the norm mismatch of the two modalities (vision and text) causing attention imbalance.
+Then we develop a compact interpretability toolkit and two mechanism-aligned interventions.
+Our toolkit comprises: (i) a Position Sensitivity Index (PSI) tool that quantifies order use, (ii) a Cross-Modality Balance (CMB) tool to analyze head-level vision vs.
+text attention with prompt isolation, and (iii) a RoPE sensitivity probe tool that measures intrinsic response to positional phase shifts.
+We also introduce a controlled synthetic benchmark (2DS) to better measure these three indicators.
+We then apply two minimal interventions to test our hypothesis: Normalize RMS scale-matching of vision embeddings post-projector, and Normalize+Multilayer combination of intermediate encoder layers.
+We do an extensive comparison with our interpretability tools to understand how our interventions can make the model behave differently.
+Our aim is understanding and measurement, instead of advancing SOTA.
+Contributions.
+(1) Diagnosis: we identify and validate a concrete mechanism —embedding-norm skew— that links modality scale to diminished positional sensitivity in the LLM.
+(2) Toolkit & data: PSI, CMB, and RoPE probe, together with a spatially focused dataset 2DS offer complementary views of spatial use.
+(3) Mechanism-aligned probes: we demonstrate that norm-induced problem is the causal mechanism behind positional insensitivity and that intermediate layers of the vision encoder can contribute more to spatial understanding.
 2 Related Work
-Vision-Language Models Recent Vision-Language Models (VLMs) combine pretrained vision encoders and large language models (LLMs) to perform image understanding tasks, such as image captioning and visual question answering [23, 24, 20, 1, 41, 28, 7, 15].
-Many approaches leverage pretrained visual encoders, notably CLIP [27], and subsequently introduce an adapter module to integrate vision embeddings into pretrained LLMs, such as Vicuna [6], enabling coherent textual outputs [23, 41].
-While these methods achieve impressive performance in descriptive tasks, they often struggle with spatial reasoning [33, 38].
-To address spatial limitations, recent efforts have focused on augmenting training datasets with extensive spatial annotations [32, 3, 5, 34].
+Vision-Language Models Recent Vision-Language Models (VLMs) combine pretrained vision encoders and large language models (LLMs) to perform image understanding tasks, such as image captioning and visual question answering (Liu et al., 2023; 2024; Li et al., 2023a; Alayrac et al., 2022; Zhu et al., 2024; Shao et al., 2023; Ding et al., 2022; Hu et al., 2022).
+Many approaches leverage pretrained visual encoders, notably CLIP (Radford et al., 2021), and subsequently introduce an adapter module to integrate vision embeddings into pretrained LLMs, such as Vicuna (Chiang et al., 2023), enabling coherent textual outputs (Liu et al., 2023; Zhu et al., 2024).
+While these methods achieve impressive performance in descriptive tasks, they often struggle with spatial reasoning (Tong et al., 2024b; Yuksekgonul et al., 2023).
+To address spatial limitations, recent efforts have focused on augmenting training datasets with extensive spatial annotations (Tong et al., 2024a; Chen et al., 2024a; Cheng et al., 2024; Trabelsi et al., 2025).
 In contrast, our approach does not rely on additional spatial labels or modifications to the original training data; instead, we introduce representation-level interventions, including embedding norm normalization and mid-layer feature extraction, which demonstrate significant improvements in VLMs’ spatial reasoning capabilities.
@@
 Neo et al.
-[26] show that vision tokens in LLaVA [23] have strong semantic properties.
-Concurrently, multiple studies have highlighted the consistent failures of VLMs to comprehend simple spatial relations such as “above,” “below,” “left,” and “right” [38, 19, 33, 4].
+(2025) show that vision tokens in LLaVA (Liu et al., 2023) have strong semantic properties.
+Concurrently, multiple studies have highlighted the consistent failures of VLMs to comprehend simple spatial relations such as “above,” “below,” “left,” and “right” (Yuksekgonul et al., 2023; Kamath et al., 2023; Tong et al., 2024b; Chen et al., 2025).
 These findings underscore a critical need for enhanced interpretability and robust spatial grounding in VLM architectures.
-Our work investigates the underlying reasons why spatial information is overshadowed in LLM, offering minimal yet effective interventions that notably strengthen the spatial reasoning ability.
-Position Encoding in Transformers Positional encoding is essential in Transformer architectures to prevent degenerating into bag-of-words representations and to effectively model sequence orders [9, 35].
-Initial efforts employed absolute positional encodings, later evolving into relative positional encodings [29] and, more recently, advanced mechanisms such as Rotary Positional Embeddings (RoPE) to better capture complex sequence dependencies [30].
-Positional encoding has proven essential for modeling both language and visual data in Transformers [8].
-Despite these developments, the specific impact of positional encodings within VLMs—where vision tokens interact with textual tokens in an LLM decoder—remains underexplored.
-Recent work [37] attributed hallucination issues to RoPE’s long-term decay property.
-Complementing these insights, our work empirically and theoretically illustrates how disproportionately large embedding norms of vision tokens can overshadow positional cues, hindering spatial reasoning even when advanced positional methods like RoPE are employed.
-3 Diagnosing the Issue: Bag of Tokens
+Our work investigates the underlying reasons why spatial information is overshadowed in LLMs, offering minimal yet effective interventions that notably strengthen the spatial reasoning ability.
+3 Motivation
+Table 1: Impact of random vision token permutation on spatial reasoning accuracy (%).
 Dataset Original Permutation Difference
@@
 CV-Bench 2D 56.59 56.26 -0.33
-Table 1: Impact of random vision token permutation on spatial reasoning accuracy (%).
 VLMs demonstrate excellent capabilities in recognizing objects within images (ventral tasks) yet consistently struggle with spatial reasoning (dorsal tasks).
-This disparity suggests that VLMs might neglect spatial cues in their internal representations.
-Open-source VLMs such as LLaVA [23] typically fuse vision embeddings with textual embeddings in transformer-based language models, relying critically on positional embeddings to encode token order.
-We hypothesize that in multimodal contexts, positional embeddings might become ineffective or suppressed, causing a loss of meaningful spatial structure.
-To investigate this, we perform two diagnostic interpretability experiments: (1) Token Permutation Test: If positional order significantly contributes to spatial reasoning, randomizing token order should substantially degrade model performance.
-We want to know how much the vision token order matters.
-(2) Spatial Compression Study: If local spatial information is crucial, drastically reducing the number of vision tokens should cause significant performance deterioration.
-We want to know how much VLM really uses local spatial information.
-Our experiments indicate that VLMs remain largely robust to these two perturbations, reinforcing the hypothesis of a “bag-of-tokens” behavior.
-3.1 Token Permutation Test
-In standard NLP transformers, positional embeddings are essential for maintaining token order.
-Similarly, if VLMs depend heavily on positional embeddings for spatial reasoning, randomizing vision token order should substantially impair spatial reasoning performance.
-We hypothesize that if token order carries meaningful spatial information, random shuffling should severely degrade performance.
-Experiment Setup: We use LLaVA 1.5 7B from [23, 24] to conduct the test.
-We randomly permute vision token embeddings obtained from the vision encoder and projection layers before feeding them into the LLM, and the performance is then compared against the original unpermuted baseline.
-Figure 1 illustrates the experiment pipeline.
-We evaluate performance on datasets including VQAv2[11], POPE [21], GQA [16], and CV-Bench[32].
-Result: As shown in  Table 1, random token permutation leads only to minor performance drops (ranging from 0.2% to 2.74%).
-This surprisingly small decline reveals that VLMs exhibit significant order-invariance, suggesting a bag of tokens tendency of VLMs.
-This is summarized as the following “Intepretabality Insight”:
-Interpretability Insight No. 1:
-VLMs exhibit limited sensitivity to vision token order, reflecting a strong “bag-of-tokens” tendency.
+Open-source VLMs such as LLaVA (Liu et al., 2023) typically consist of a vision encoder and an LLM.
+One central question we want to know is: How do VLMs use the spatial information?
+Conceptually, there are two sources for the spatial information: vision encoder and LLM position embedding.
+The vision encoder encodes spatial information within vision tokens.
+LLM position embedding directly represents the spatial information of the input tokens’ order.
+We are particularly interested in how much the LLM leverages position embedding for spatial information.
+To understand the positional sensitivity, we use LLaVA 1.5 7B (Liu et al., 2023; 2024) to randomly permute vision token embeddings after processing by the vision encoder and the MLP projection layers, but right before they are fed into the LLM (Figure 1).
+It effectively changes the position embedding for each token.
+The performance is then compared against the original unpermuted baseline.
+We evaluate performance on datasets including VQAv2 (Goyal et al., 2017), POPE (Li et al., 2023c), GQA (Hudson & Manning, 2019), and CV-Bench (Tong et al., 2024a).
+As shown in Table  1, random token permutation leads only to minor performance drops (ranging from 0.2% to 2.74%).
+This surprisingly small decline reveals that current VLMs evaluated with the corresponding benchmarks are largely insensitive to positional information.
 Figure 2: Performance impact of vision token compression on standard benchmarks (GQA, CV-Bench 2D, and POPE).
 Only minor accuracy degradation occurs, even under extreme token compression (down to a single token).
-3.2 Spatial Compression Study
-Recent work showed VLMs focus on object semantic tokens[26].
-To further probe the use of spatial information by VLMs, we progressively compress the vision tokens, effectively removing fine-grained spatial details including detailed positional and geometric relationships among image patches.
-If spatial details are crucial for evaluation, performance should sharply decline as spatial resolution decreases.
-Conversely, stable performance despite extreme compression may indicate either that the model relies predominantly on semantic features or that existing benchmarks do not effectively challenge spatial reasoning capabilities.
-Experiment Setup: Using the LLaVA 1.5 7B model, we systematically reduce the number of vision tokens from the original 576 down to just 1.
-Models were retrained for token counts (256, 64, 16, and 1) using average pooling after the MLP projection layer.
-This method severely limits local spatial details while preserving some semantic content.
-Performance impacts are evaluated across various benchmarks.
-Result: As illustrated in  Figure 2, reducing the vision tokens from 576 down to just 1 for standard benchmarks results in a minor accuracy drop (-8.5% the worst case).
-Despite losing approximately 99.8% of spatial resolution, the model’s performance remains robust.
-This outcome suggests two interrelated explanations: first, VLMs may primarily rely on semantic cues rather than local spatial details when solving typical benchmark tasks; second, the current benchmarks themselves might not effectively test the model’s capacity for detailed spatial reasoning.
-The limited sensitivity of these benchmarks to extreme spatial compression could reflect both model behavior and benchmark limitations.
-This leads us to the following insight:
-Interpretability Insight No. 2:
-For current benchmarks, VLMs can predominantly rely on “ventral” semantics rather than “dorsal” spatial cues to have a decent performance.
-4 Why Does This Happen? Analysis of Embedding Norm Suppression
-In earlier experiments, we observed that spatial information in VLMs tends to be minimal or overshadowed by strong global semantic signals.
-Here, we further investigate the underlying cause by analyzing embedding norms, specifically examining how large vision embedding magnitudes overshadow positional encodings within attention mechanisms.
-4.1 Embedding Norm Analysis
-Motivation Positional embeddings were initially introduced into Transformer architectures to maintain token order and avoid the “bag-of-words” effect [35].
-While VLMs also utilize positional encodings during training, we hypothesize these encodings become ineffective if vision embeddings possess disproportionately large magnitudes, thus dominating positional signals during attention computations.
-To investigate this hypothesis, we conduct an empirical embedding analysis using the COCO validation dataset (5000 image-text pairs) [22].
-Specifically, we extract vision and text embeddings after MLP projection but before input into the language model and compute the L2 norms for each token type.
-Figure 3 illustrates the embedding norm distributions.
-Figure 3: Distribution of L2 norms for vision and text tokens in COCO validation dataset (log scale).
+Another way to investigate the role of position is to eliminate the position embedding completely.
+Recent work showed VLMs focus on object semantic tokens (Neo et al., 2025), and we want to know how much spatial information is encoded into the vision tokens and how much is the result of the LLM position embedding.
+We progressively compress the vision tokens, effectively removing the position embedding’s effect on the spatial information, including detailed positional and geometric relationships among image patches.
+We aim to push it to the extreme to see how much information is relied on the LLM position embedding or if they are already encoded in the vision token.
+Using the LLaVA 1.5 7B model, we retrained with vision token counts of 256, 64, 16, and 1 (originally 576), using average pooling after the MLP projection layer.
+As illustrated in Figure  2, reducing the vision tokens from 576 down to just 1 for standard benchmarks results in a minor accuracy drop (-8.5% the worst case).
+Despite losing approximately 99.8% (from 576 to 1) of spatial resolution, the model’s performance remains robust.
+Concurrent work also reached similar conclusion and showed robustness of compression (Chai et al., 2025).
+This outcome suggests two possible reasons.
+First, VLMs may primarily rely on vision encoder’s spatial encoding, and position embedding in the LLM has low effects.
+Second, the current benchmarks might not effectively test the model’s capacity with the LLM part of the spatial information encoding.
+4 Why Does This Happen? Embedding Norm Suppression
+The previous section showed a striking failure of positional encoding: permuting vision tokens often leaves predictions minor change, i.e., an unintended “bag-of-words” behavior despite the presence of positional encodings (Vaswani et al., 2017; Yuksekgonul et al., 2022).
+To probe why, we conduct an empirical norm analysis on COCO validation (5k image–text pairs) (Lin et al., 2015).
+We extract vision embeddings after the MLP projector and before the language model, compute their $L_{2}$ norms, and compare to text-token norms.
+Figure 3 shows the distribution comparison.
+Figure 3: Distribution of $L_{2}$ norms for vision and text tokens in COCO validation dataset (log scale).
 Vision token norms range between $10^{1}$ and $10^{3}$, while text token norms range between $3\times 10^{-1}$ and $10^{0}$.
-As shown in Figure 3, vision embeddings consistently have 1-2 orders of magnitudes greater than text embeddings but can reach up to 3 orders of magnitude greater at its upper range.
-Although directional alignment (e.g., cosine similarity) typically receives primary focus in multimodal alignment due to inner product [14, 17, 25], such large magnitude discrepancies may suppress positional signals significantly.
-4.2 Theoretical Explanation
-We provide a concise theoretical justification of how differences in embedding norms significantly impact the effectiveness of positional encodings in Transformer attention mechanisms.
-A detailed mathematical derivation is presented in the Appendix A but here is a summary.
-Summary of Proof (detailed derivation in Appendix A).
-Consider a Transformer self-attention block employing Rotary Positional Embeddings (RoPE) [31].
-RoPE applies orthonormal rotations to query and key vectors with dimension $d$ ($\mathbf{q}_{i},\mathbf{k}_{i}\in\mathbb{R}^{d}$), preserving their norms.
-After RoPE transformation, we denote these vectors as $\mathbf{q}_{i}^{\prime},\mathbf{k}_{i}^{\prime}\in\mathbb{R}^{d}$.
-The attention logits between token pairs are computed via the scaled dot-product:
-$\mathrm{logits}_{ij}=\frac{\mathbf{q}_{i}^{\prime}\cdot\mathbf{k}_{j}^{\prime}}{\sqrt{d}}=\frac{\|\mathbf{q}_{i}\|\|\mathbf{k}_{j}\|\cos{\phi}}{\sqrt{d}}$ (1)
-where $\phi$ denotes the positional angle difference.
-From this expression, we observe directly that larger embedding norms increase the scale of logits.
-When vision embedding norms substantially exceed text embedding norms (often by an order of magnitude), vision-to-text logits dominate softmax computations, resulting in disproportionately high attention scores.
-To maintain coherent textual outputs aligned with the language model’s training, the model implicitly learns to downscale these vision-related logits.
-This downscaling leads to empirically observed low attention weights for vision tokens, consistent with findings in recent studies [4].
-Consequently, the positional sensitivity explicitly encoded by RoPE diminishes substantially, as lower attention weights inherently reduce the positional gradient (further detailed in Appendix A).
-Interpretability Insight No. 3 The practical effectiveness of positional encoding mechanisms’ (such as RoPE) is severely limited by embedding magnitude discrepancies between modalities.
-Normalizing or rescaling embedding norms thus emerges as a crucial intervention, restoring positional sensitivity and enabling positional embeddings to meaningfully influence attention computations.
-This insight extends broadly to multi-modal alignment scenarios: effective alignment between pretrained models should not rely solely on embedding similarities (such as cosine similarity) but must also carefully consider embedding norm differences.
-Otherwise, positional information can become suppressed, leading to minimal influence on model predictions.
-5 Restoring Spatial Awareness: A Controlled Study
-Our previous analyses indicated that spatial cues in VLMs are often overshadowed by strong semantic signals and disproportionately large embedding norms.
-To validate these findings and explore methods for enhancing spatial representations, we designed a controlled study employing two interpretability-informed strategies:
-  1. 1.
-    Synthetic Dataset: A carefully constructed two-dimensional (2D) spatial benchmark designed explicitly to eliminate semantic shortcuts, ensuring that spatial relationships alone drive task success.
-  2. 2.
-    Interpretability-Informed Model Adjustments: Targeted architectural modifications aimed at balancing positional cues and mitigating the overwhelming dominance of semantic embeddings.
-Combining these controlled experimental conditions with interpretability-driven adjustments allows us to investigate the feasibility of reviving effective spatial reasoning—analogous to restoring a functional “dorsal stream”—in VLMs.
+In LLaVA 1.5 model, vision embeddings are typically 1–2 orders of magnitude larger than text embeddings, with an upper tail approaching 3 orders.
+While alignment work often emphasizes direction (cosine similarity) because logits are inner products (Hessel et al., 2021; Huh et al., 2024; Maniparambil et al., 2024), such large magnitude gaps can dominate the scaled dot-product and suppress positional signals.
+Our analysis focuses on the LLaVA architecture, which uses a LLaMA decoder (Grattafiori et al., 2024).
+LLaMA employs a pre-Norm design where each sublayer follows a specific pattern: (1) the input hidden states is first normalized by RMSNorm (Zhang & Sennrich, 2019b), (2) the computational block (e.g., self-attention) processes this normalized input to produce an update, and (3) this update is added back to the original, unnormalized residual.
+RoPE is applied within the self-attention block to the normalized inputs (Su et al., 2024).
+We hypothesize that the large vision norms in the residual path are not normalized.
+It reduces the RoPE effectiveness applied in the attention when we consider the magnitude difference between attention outputs and unnormalized vision tokens.
+It might also let the vision token direction dominate due to small changes to the residual in early layers, which could cause vision attention bias.
+[…]
+If the above mentioned hypothesis is correct, then the hidden states of vision tokens should be much larger than text tokens in early layers.
+On COCO val (5k), we partition tokens into vision ($V$) and user text ($T$).
+We exclude the fixed system prompt for clarity.
+We compute each token’s $L_{2}$ norms and average over respective tokens.
+Figure 4 shows the results.
+Vision hidden states are an order of magnitude larger than text in early layers, and the disparity persists until roughly mid‑depth (Figure 4 right).
+This directly supports the residual‑scale suppression account.
+More detailed explanation is given in Appendix A.
+Overall we want to measure the RoPE effectiveness directly to see its impact.
+To measure the RoPE effect directly, we can check the partial derivative of the attention w.r.t.
+position phase ($\phi$)’s changes.
+Let logits be $\ell_{j}=\langle q^{\prime},k^{\prime}_{j}\rangle/\sqrt{d}$ and $\alpha_{k}=\mathrm{softmax}(\ell)_{k}$ be attention weights, and partition tokens into vision ($V$) and text ($T$).
+Define the group masses $\alpha_{V}=\sum_{v\in V}\alpha_{v}$, $\alpha_{T}=\sum_{t\in T}\alpha_{t}$, and the group‑average logit derivatives
+$g_{V}\;=\;\sum_{v\in V}\frac{\alpha_{v}}{\alpha_{V}}\,\frac{\partial\ell_{v}}{\partial\phi},\qquad g_{T}\;=\;\sum_{t\in T}\frac{\alpha_{t}}{\alpha_{T}}\,\frac{\partial\ell_{t}}{\partial\phi}.$
+Then (derivation in Appendix A) the change of total vision attention mass w.r.t. the positional phase $\phi$ is
+$\frac{\partial\alpha_{V}}{\partial\phi}\;=\;\alpha_{V}\,\alpha_{T}\,\bigl(g_{V}-g_{T}\bigr).$ (1)
+This derivation establishes a direct link between the shift in vision attention mass ($\frac{\partial\alpha_{V}}{\partial\phi}$) and the underlying positional sensitivity of the vision logits ($g_{V}$).
+According to this relationship, a very small attention shift would mean low positional sensitivity.
+Conversely, a large attention shift would indicate high positional sensitivity, suggesting that RoPE is functioning effectively.
+This principle provides the mathematical support for RoPE sensitivity probe in the next section.
+5 Interpretability Toolkit: Understand who is looking at what
+We develop a compact toolkit to probe where spatial information is (or is not) used inside VLMs.
+The tools target three complementary questions: (i) Does position embedding matter?
+(Position Sensitivity), (ii) Which components actually look at vision?
+(Cross-Modality Balance at the head level), and (iii) How sensitive is the model to positional phase shifts?
+(RoPE Sensitivity).
+5.1 Position Sensitivity Index (PSI)
+As discussed in Section 3, shuffling vision tokens often leaves answers mostly unchanged, suggesting limited use of spatial order through position embedding.
+We quantify this with the Position Sensitivity Index:
+$\mathrm{PSI}\;=\;\frac{\mathrm{Acc}(\text{original order})-\mathrm{Acc}(\text{permuted order})}{\mathrm{Acc}(\text{original order})}.$ (2)
+PSI is simple, scale-free, and interpretable: $\mathrm{PSI}=0$ means complete order invariance, and larger values indicate greater reliance on the token order.
+We report PSI at both the model and dataset levels since it could be both sides’ issues.
+This lets us compare models and datasets on a common measure.
+5.2 Cross Modality Balance (CMB)
+Functional behavior in Transformers is often head-specific, for example there are induction, copy, memory, summary, truthfulness attention heads in LLM (Zheng et al., 2024; Bietti et al., 2023; Tigges et al., 2023; Olsson et al., 2022; Li et al., 2023b).
+We therefore work at head granularity to give a closer look at the vision and text modality balance.
+Let $V$ and $T$ denote the index sets of vision and text tokens, respectively. For a fixed decoding step and attention head $h$ with attention weights $a_{\cdot,h}$, we have:
+$\mathrm{CMB}_{h}\;=\;\frac{\sum_{v\in V}a_{v,h}}{\sum_{i\in V\cup T}a_{i,h}}.$ (3)
+where $\mathrm{CMB}_{h}\in[0,1]$ measures how much attention head $h$ attends to vision versus text.
+In practice, we summarize the CMB with layer-by-layer head heatmaps.
+We can visualize which head is more focused on the vision and which is more focused on the text.
+5.3 RoPE Sensitivity Probe
+To empirically quantify the practical impact of RoPE’s positional information on the models’ attention mechanism, we introduce a RoPE sensitivity probe to measure how much positional phase (RoPE) affects what the model predicts.
+For a fixed query position, we first extract the query vector and all key vectors to compute the baseline attention score, denoted as $\alpha$.
+We then create a perturbed state by applying an additional RoPE rotation of $\Delta$ steps exclusively to vision keys.
+The query vector and text key vectors remain unchanged and preserve original causal masking.
+Finally, we recompute the attention score, $\alpha_{\Delta}$, using the perturbed vision keys.
+This perturbation allows us to isolate how the positions of visual evidence shift relative to a fixed linguistic query.
+From this procedure, we report two key measures that serve as a finite-difference approximation of terms in Equation 4, which is also listed here for its importance:
+$\frac{\partial\alpha_{V}}{\partial\phi}\;=\;\alpha_{V}\,\alpha_{T}\,\bigl(g_{V}-g_{T}\bigr).$ (4)
+(A) Attention-level sensitivity.
+As an approximation of $\frac{\partial\alpha_{V}}{\partial\phi}$, $\Delta\alpha_{V}$ is defined as the finite difference between $\alpha$ and $\alpha_{\Delta}$ with step $\Delta$.
+This captures the observed change in how much attention mass moves among vision tokens under a RoPE phase shift.
+(B) Intrinsic (logit-level) sensitivity.
+We can also directly measure $\Delta g_{V}$ which is the group-average logit derivative w.r.t.
+the RoPE angle.
+It is an approximation of the $g_{V}$ term.
+In experiments we report both the attention-level difference and the normalized logit-level difference.
 5.4 A 2D Synthetic Spatial Benchmark (2DS)
-Standard image and text benchmarks often unintentionally provide semantic or contextual shortcuts, enabling models to leverage object co-occurrence or linguistic biases without genuinely understanding spatial relationships.
-We particularly hope to respond to the potential inadequacies in current benchmarks from our spatial compression study (Section 2).
-To isolate and rigorously evaluate spatial reasoning, we limit confounding factors such as real-world object frequencies and linguistic priors.
-Our benchmark specifically targets 2D spatial relationships directly influenced by positional embeddings.
+Figure 5: 2DS examples. Left: two-object layouts; right: three/four-object layouts. Semantics are simple to focus on spatial relations.
+Standard benchmarks often admit semantic shortcuts (object co-occurrence, language priors), masking spatial failure modes (Section 3 and Fig.
+2).
+Our dataset 2DS removes these shortcuts: scenes contain colored shapes placed at controlled locations; queries require absolute (e.g., “bottom”) or relative (e.g., “left of”) relations, with semantics (color/shape) serving only as identifiers.
+We generate multiple meta-categories by number of objects (2–6), randomize positions within each, and enumerate questions across {color, shape, color+shape} $\times$ {absolute, relative}.
+Full construction details are in Appendix B.
+We use the same methods from Section 3 to test 2DS’s position effectiveness, and the results are reported in Appendix B.2.
+It shows a much higher sensitivity over position compared to other datasets.
+6 Probing Spatial Awareness: A Controlled Study
+Guided by the diagnostics above, we test two concrete hypotheses about why VLMs underuse spatial cues under perturbations:
+H1 (Norm-induced position embedding weakness). A modality norm mismatch (vision $\gg$ text) inflates vision token norms and attenuates RoPE effects.
+H2 (Mid-layer spatial richness). Spatial detail is concentrated in intermediate vision-encoder layers, and final layers are more semantic, less spatial (Wang et al., 2023; Jiang et al., 2024).
+6.1 Interventions and Setup
+We use LLaVA-1.5 (7B) (Liu et al., 2023) as the baseline, reproducing its two-stage training and data.
+We propose and evaluate the following two minimal interventions.
+From the derivative in equation 4, H1 predicts that reducing the vision norm can improve positional sensitivity.
+H2 predicts intermediate layers can add additional gains from richer spatial evidence.
++ Normalize (H1).
+RMS normalization (Zhang & Sennrich, 2019a) of vision embeddings after the linear projector and before feeding the LLM, calibrated to typical text-token norms (empirically: mean $\approx 0.83$, max $\approx 1.22$).
++ Normalize + Multilayer (H2).
+Concatenate intermediate vision features to the vision token dimension (layers 12, 16, 20, 24) following (Jiang et al., 2024), then apply the same normalization.
+This aims to inject fine-grained spatial signals while controlling scale.
+Multilayer adds a small compute overhead in theory with extra embedding expanded from 1024 to 4096 dimension.
+We showed empirically it is negligible in Appendix D.
+6.2 Position Sensitivity Index
+Table 2: Position Sensitivity Index (%).
+Data LLaVA 1.5 + Normalize + Normalize + Multilayer
+VQAv2 1.09 2.03 2.08
+POPE 0.34 0.35 0.46
+GQA 4.47 6.10 5.64
+CV-Bench 2D 0.57 11.31 9.72
+2DS 41.07 61.21 54.21
+PSI could measure a two-fold problem – model and dataset.
+Table 2 compares both for PSI.
+If we fix one dataset, we can compare across the models, and vice versa.
+Given a baseline model, PSI isolates model’s order use: if shuffling vision tokens barely moves accuracy compared to the baseline, the model likely underutilizes spatial arrangement.
+We can also compare it across benchmarks to see which one is more sensitive to the position changes.
+Particularly, we find 2DS has a significantly higher PSI compared to other datasets across different models, which means that 2DS can test position embedding more effectively than existing benchmarks.
+(1) Normalization substantially increases PSI across diverse benchmarks, consistent with our hypothesis H1: reducing norm skew restores RoPE leverage and raises order-use.
+(2) Multilayer occasionally trails normalization on PSI (e.g., CV-Bench 2D, 2DS), yet as shown in Table 3 and Appendix E it can yield higher accuracy, indicating that models can exploit spatial content from vision encoder features even if their explicit sensitivity to token permutations is moderate.
+In short, PSI probes order use, while accuracy reflects spatial competence.
+Table 3: 2DS accuracy (%) by attribute/relation. $\uparrow$ higher is better. $\Delta$ vs. baseline in parentheses.
+Category LLaVA 1.5 + Normalize + Normalize + Multilayer
+Color_abs. $\uparrow$ 79.60 88.00 (+8.40) 87.80 (+8.20)
+Color_rel. $\uparrow$ 37.00 42.40 (+5.40) 47.20 (+10.20)
+Shape_abs. $\uparrow$ 78.60 79.00 (+0.40) 82.20 (+3.60)
+Shape_rel. $\uparrow$ 34.40 32.80 (-1.60) 40.80 (+6.40)
+Shape_color_abs. $\uparrow$ 70.80 71.80 (+1.00) 80.20 (+9.40)
+Shape_color_rel. $\uparrow$ 39.40 41.80 (+2.40) 50.60 (+11.20)
+Overall Acc. $\uparrow$ 56.63 59.30 (+2.67) 64.80 (+8.17)
+6.2.1 Which attributes are important?
+Table 3 showed that +Normalize chiefly helps categories where color anchors spatial relations (Color_abs./rel., Shape_color_*), suggesting that final-layer features retain strong color cues but weaker geometry.
+By contrast, +Normalize+Multilayer substantially lifts relative categories across the board, aligning with H2: intermediate features furnish finer geometry and local layout that compose with color cues.
+The largest jumps appear where both feature identity and geometry must be integrated.
+On 2DS, +Normalize attains the highest PSI (61.21%) while +Normalize+Multilayer yields the highest accuracy (64.8%).
+This suggests that restoring positional leverage (H1) and supplying spatial content (H2) are complementary spatial information encoding.
+Their relative contributions differ by metric.
+6.3 Cross Modality Balance: Where do attention heads look?
+System prompt confound.
+Prior work reports very low vision attention vs.
+text (Chen et al., 2024b; 2025).
+Our toolkit separates system prompt from user text and vision, revealing why.
+Text can break into two categories, the system prompt, and the user input.
+We compare those two and the vision attentions.
+As in Fig.
+6, the static system prompt absorbs a large, image-invariant portion of attention (72.1%), masking vision use.
+It might act as an attention sink (Xiao et al., 2024).
+It provides a new explanation on why prior work reports very low vision attentions.
+When we exclude the system prompt from our analysis, a different picture emerges: in the original VLMs, the attention mass on vision tokens is significant higher than the user input text.
+However, we believe the apparent visual dominance is not a sign of a healthy mechanism.
+Instead, it is a ”brute force” consequence of the large vision token norms, as described in Section 4.
+The normalization can balance the vision and user input text attention as we show in Appendix F.
+Figure 6: LLaVA 1.5 7B attention share average over COCO Validation dataset.
+Green is system prompt, orange is vision token, and blue is user input text.
+Left graph includes system prompt.
+Right graph just compares vision and input text.
+Modality-integration patterns.
+To understand how the VLMs integrate the vision and text modality, we visualize the CMB with heatmaps across the layers and attention heads.
+It can give us a granular sense of what each attention head focuses on, vision or user input text.
+Figure 7 showed the CMB heatmap visualization average across 5k COCO val results.
+With prompt excised for clarity, LLaVA-1.5 shows many vision-specialized heads scattered across depth.
+It aligns with the analysis in Section 4 that the vision tokens could dominate the attention calculation.
++Normalize exhibits a vision to text progression: early layers lean on vision, deeper layers shift toward text, indicating more balanced cross-modal exchange once norm disparity is removed.
++Normalize+Multilayer displays a striking stage-like pattern: the first two layers emphasize vision, an early-block pivot concentrates on text, and deeper layers re-amplify vision with smooth transition.
+We view it as a pipeline that extract, align, then integrate.
+We also provide more individual image visualization and more dettails of calculation in Appendix G.
+[…]
+The performance of models over standard benchmarks in Appendix E show that +Normalize+Multilayer is the best, then +Normalize, then original.
+It suggest original model’s “visual dominance” does not mean a better visual performance.
+With large vision activations in the residual stream (Sec.
+4), the normalized state is pre‑aligned to the vision subspace, so queries favor vision keys by default, and high CMB reflects a directional bias towards vision, not use question conditioned retrieval.
+As our attention visualizations and entropy measures indicate (Appx.
+H), baseline vision maps are diffuse and often focus on the wrong target objects asked in text.
+In contrast, +Normalize+Multilayer’s staged vision-text-vision has the most focused (lowest‑entropy) attention maps and the best spatial accuracy.
+Figure 8: RoPE sensitivity (higher = larger change under phase shift) across layers.
+6.4 RoPE sensitivity: How does normalization help?
+One key question we want to understand is whether the normalization helped or not.
+We already see that the normalization can effectively balance the hidden states in Section 4 Figure 4.
+However, to have a more indepth understanding, we want to know how RoPE directly changes vision.
+Our RoPE probe perturbs positional phase for vision keys only and measures (Fig.
+8): (i) attention-level change $\Delta\alpha$, and (ii) the normalized logit change $\Delta g_{V}$, which factors out attention imbalance as we observed from CMB results.
+We show the RoPE probe with $\Delta=1$ applied to the vision keys, and we report the RoPE sensitivity measures over COCO dataset average in Figure 8.
+We see both $\Delta\alpha_{V}$ and $\Delta g_{V}$ increase the sensitivity in early 0-5 layers over RoPE changes on vision tokens.
+Both +Normalize and +Normalize+Multilayer show markedly higher sensitivity than baseline in early layers (0–5).
+(Fig.
+8).
+This matches H1: reducing norm skew raises sensitivity of the RoPE effects.
+We believe it shows that normalization helps restoring vision token positional information from RoPE, especially in early layers.
+Position embedding increase the spatial channel information, improving both order-use (PSI) and task competence (2DS).
+7 Discussion and Future Work
+What does this reveal about VLM internal mechanisms?
+Spatial signals enter through (i) vision-encoder features and (ii) the LLM’s positional mechanism (RoPE).
+Our probes indicate that both are necessary but they play different roles.
+Scale matching (Sec.
+4) restores order sensitivity (higher PSI and RoPE probes).
+Exposing intermediate vision layers supplies finer geometry and local layout that raise spatial competence even when PSI changes modestly.
+The division of labor therefore shifts with design: +Normalize leans on RoPE-driven order, and let the LLM do more works on spatial information.
++Normalize+Multilayer delegates more spatial work to the vision encoder.
+Why are raw vision attentions “low”?
+Layer-averaged attention under-counts visual use because the static system prompt behaves like an attention sink.
+Once we excluded it for analysis only, many attention heads are vision-preferring (Section 6.3), especially early/mid layers.
+Normalization rebalanced this imbalance caused by residual norm (vision share drops from $\sim$75% to $\sim$54% as shown in Appendix F).
+Results suggest a more balanced vision and text could perform better, possibly due to a better question conditioned retrieval.
+Vision attention are also less diffused and focus less on the wrong object as shown in Appendix H.
+[…]
+Limitations and Implications Our normalization is blunt and can reduce dynamic range, especially for 3D cues.
+It reduces the range of the expression and might lead to reduced 3D performance.
+Future work should consider a more sophisticated normalization alignment.
+Our work is limited to the LLaVA style model, which has a pre-trained vision encoder align with a pre-trained LLM through a projection layer.
+Recent VLMs train vision encoders from scratch and might not have the problem we mentioned (Bai et al., 2025; Grattafiori et al., 2024).
+8 Conclusion
+We show how VLMs use spatial information, and we provide a compact interpretability toolkit to make this visible and measurable.
+Best to our knowledge, we are the first to connect embedding–norm skew to diminished positional sensitivity in the LLM (RoPE), with analysis and experiments supporting the link.
+This mechanistic understanding reveals how modality-specific scales can inadvertently suppress cross-modal information flow.
+We introduce four simple, targeted tools for spatial information measurements: PSI for order use, CMB for head-level modality balance with prompt isolation, a RoPE sensitivity probe for intrinsic phase response, and 2DS, a controlled spatial benchmark.
+We showed two minimal changes and clarified design levers: Normalize (scale-match vision to text) restores RoPE leverage and raises PSI and RoPE sensitivity; Multilayer adds mid-layer geometry and boosts performance accuracy.
+We believe those tools and insights are valuable for future VLMs design considerations.
+Acknowledgments
+The work is supported by the National Science Foundation (NSF) through Awards #2131186 (CISE-MSI), #1827505 (PFI), and the US Air Force Office of Scientific Research (AFOSR) via Award #FA9550-21-1-0082.
+The work is also supported by a College-wide Research Vision (CRV) Fund from the CCNY Provost’s Office and the Google CyberNYC Initiative.
+This work used Google Cloud through the CloudBank project, which is supported by National Science Foundation Award #1925001.
+We specifically thank Dr.
+Ping Ji’s generous support for computing resources.
+Appendix A Theoretical Analysis of Large-Norm Effects on RoPE
+We consider a pre-norm Transformer block (as used in LLaVA-style decoders). For token state $h^{(l)}\!\in\!\mathbb{R}^{d}$ at layer $l$,
+$\underbrace{a^{(l)}(\phi)}_{\text{attn out}}\;=\;\mathrm{Attn}\!\big(\mathrm{RMSNorm}(h^{(l)}),\phi\big),\qquad h^{(l+1)}\;=\;\underbrace{r^{(l)}}_{\text{residual carry}}\;+\;a^{(l)}(\phi),$
+where $\phi$ denotes the RoPE phase and $r^{(l)}\!=\!h^{(l)}$ (the pre-attention input).
+Keys and queries are formed from per-token RMS-normalized inputs; RoPE applies a norm-preserving rotation to $q,k$ (Su et al., 2023).
+However, the residual is added to every normalization results even after the post layer norm.
+It causes the unbalanced vision token norm to keep passing on.
+Residual-scale suppression (directional lemma).
+With $h^{(l+1)}=r^{(l)}+a^{(l)}(\phi)$ and $u(\phi)=h^{(l+1)}/\|h^{(l+1)}\|$.
+With some quotient rule and replacing $\frac{\partial\|h^{(l+1)}\|}{\partial\phi}=\frac{(h^{(l+1)})^{T}}{\|h^{(l+1)}\|}\,\frac{\partial h^{(l+1)}}{\partial\phi}$ we can have
+$\frac{\partial u}{\partial\phi}=\frac{(I-uu^{\top})}{\|h^{(l+1)}\|}\,\frac{\partial a^{(l)}}{\partial\phi}.$
+When $\|r^{(l)}\|\!\gg\!\|a^{(l)}\|$ the magnitude of $\partial u/\partial\phi$ is suppressed by $\sim 1/\|r^{(l)}\|$ in the vision token part.
+Particularly, the current layer’s sensitivity is proportional to the attention sensitivity of the previous layer divided by a residual norm.
+A large vision token norm would mean much less effective position sensitivity.
+One of the key terms is the attention level sensitivity.
+It complements attention-level analyses by explaining why even nontrivial attention shifts may yield small representational changes when the residual dominates.
+B. Attention-level phase sensitivity (group form)
+Within a given head, let logits be $\ell_{j}=\langle q^{\prime},k^{\prime}_{j}\rangle/\sqrt{d}$, and define attention weights $\alpha_{j}=\mathrm{softmax}(\ell)_{j}=\frac{exp(\ell_{j})}{\sum_{i}exp(l_{i})}$.
+Partition keys into vision $V$ and text $T$, and define group masses
+$\alpha_{V}=\sum_{v\in V}\alpha_{v},\qquad\alpha_{T}=1-\alpha_{V}.$
+Let the group-average logit derivatives be
+$g_{V}=\sum_{v\in V}\frac{\alpha_{v}}{\alpha_{V}}\frac{\partial\ell_{v}}{\partial\phi},\qquad g_{T}=\sum_{t\in T}\frac{\alpha_{t}}{\alpha_{T}}\frac{\partial\ell_{t}}{\partial\phi}.$
+A standard calculation with some quotient rule from softmax yields the single vision key given a fixed query:
+$\frac{\partial\alpha_{v}}{\partial\phi}=\alpha_{v}\left(\frac{\partial\ell_{v}}{\partial\phi}-\sum_{k}\alpha_{k}\frac{\partial\ell_{k}}{\partial\phi}\right)$ (5)
+We can use the above notation and get the group-level sensitivity by aggregating all vision attentions then substitute with $\alpha_{V},\alpha_{T},g_{V},g_{T}$
+$\frac{\partial\alpha_{V}}{\partial\phi}$ $=\alpha_{V}\left(g_{V}-\sum_{k}\alpha_{k}\frac{\partial\ell_{k}}{\partial\phi}\right)$ (6)
+$=\alpha_{V}\left(g_{V}-(\sum_{v\in V}\alpha_{v}\frac{\partial\ell_{v}}{\partial\phi}+\sum_{t\in T}\alpha_{t}\frac{\partial\ell_{t}}{\partial\phi})\right)$ (7)
+$=\alpha_{V}\left(g_{V}-(\alpha_{V}g_{V}+\alpha_{T}g_{T})\right)$ (8)
+$=\alpha_{V}\alpha_{T}\left(g_{V}-g_{T}\right)$ (9)
+Eq. equation LABEL:eq:group-derivative-appendix is an attention-level statement, and we can directly measure the change without the need to consider the previous layer.
+Small-phase probe.
+If we rotate only vision keys by a small phase $\Delta$ (queries fixed), then $g_{T}$ is unchanged and to first order
+$\Delta\alpha_{V}\;\approx\;\alpha_{V}(1-\alpha_{V})\,\Delta g_{V},$ (11)
+This factorization underlies our probe: $\Delta\alpha_{V}$ grows when (i) the balance factor $\alpha_{V}(1-\alpha_{V})$ is large (less saturation) and/or (ii) the intrinsic phase response $\Delta g_{V}$ is large.
+Empirically we can capture $q,k$ inside attention and conduct position perturbation, so we can measure $\Delta\alpha_{V}$ and $\Delta g_{V}$ with attention-level sensitivity.
+The effective positional change is modulated by the residual.
+Reducing the interface scale (our +Normalize) decreases $\|r^{(\ell)}\|$ for vision tokens in early layers, mitigating residual-scale suppression and making attention-level phase shifts consequential for subsequent computation.
+It also confirmed with our observation that +Normalize increases early layers’ position sensitivity in Section 6.4.
+Appendix B 2DS Dataset Details
 We constructed a synthetic dataset that: (1) Ensures spatial cues are essential for correctly answering queries.
@@
 These queries illustrate that object attributes (color, shape) serve primarily as identifiers, with positional relationships being crucial for correct responses.
-5.1.1 Dataset Construction Details
+B.1 Dataset Construction Details
 We created meta-categories containing images with 2, 3, 4, 5, and 6 objects.
@@
 Note the emphasis is on spatial relationship reasoning.
-Figure 4 provides representative examples, emphasizing the simplicity of semantic attributes and the centrality of spatial relationships.
-Figure 4: Illustrative scenes from our 2DS Dataset. The left two images show two object arrangements, while the right two images illustrate more complex three and four object arrangements.
-5.2 Interpretability-Informed Model Adjustments
-Our interpretability analyses identified embedding norm suppression and insufficient usage of local geometric information as critical limitations to effective spatial reasoning in VLMs.
-Guided by these insights, we propose two straightforward and interpretable adjustments: (1) Vision Token Normalization and (2) Utilizing Intermediate-Layer Features.
-Vision Token Normalization Our prior analyses revealed that vision embedding magnitudes significantly exceed those of positional embeddings, effectively suppressing positional signals during attention computations.
-To address this imbalance:
-  • Norm Matching: We measured the distribution of typical text embedding norms (mean $\approx 0.83$, max $\approx 1.22$).
-  • RMS Normalization: We applied RMS normalization [39] that is used in LLM [6] to vision encoder outputs to match them closely with textual embedding norms.
-This normalization reduces the dominance of semantic signals, allowing positional cues to emerge more clearly.
-Intermediate-Layer Features (Local Geometry) Current VLM architectures typically use highly aggregated, semantically-rich final-layer vision features from CLIP, which excel in semantic alignment but may lack fine-grained spatial details [36].
-Recent studies suggest that intermediate layers of CLIP encoders retain richer local spatial information beneficial for VLMs [18].
-Thus, we investigate whether directly incorporating earlier spatial features improves spatial reasoning.
-Specifically, we extracted features from intermediate vision encoder layers (12, 16, 20, and 24), inherently containing more detailed spatial information reminiscent of the dorsal-stream processing pathway.
-These mid-layer features explicitly encode fine-grained local spatial information, potentially crucial for accurately answering purely spatial queries in our synthetic dataset.
-However, an interpretability-informed ablation study (Section 7.2) reveals a semantic-spatial trade-off: while intermediate features significantly boost purely spatial task performance, they might also reduce the LLM’s incentive to develop explicit positional sensitivity within its attention mechanism.
-6 Experiments
-Figure 5: Accuracy comparison across varying numbers of objects. Our interpretability-informed adjustments yield consistent improvements, especially as spatial complexity increases.
-We now evaluate the effectiveness of our interpretability-informed interventions, namely vision embedding normalization and intermediate-layer feature extraction, in enhancing spatial reasoning in VLMs.
-Our experiments cover results on both a controlled synthetic spatial benchmark and several standard vision-language benchmarks.
-6.1 Experimental Setup
-We use LLaVA 1.5 (7B) [23] as our baseline model, which is chosen for its open-source availability and transparent training details.
-We replicate LLaVA 1.5’s exact two-stage training process and dataset usage for fair comparison.
-Based on the interpretability insights discussed earlier, we evaluate two proposed variants:
-  • + Normalize: Vision embeddings normalized according to the procedure detailed in Section 5.2.
-  • + Normalize + Multilayer: Normalized vision embeddings augmented with intermediate-layer features to retain fine-grained spatial information ( Section 5.2).
-6.2 Results on Synthetic Dataset
-Table 2 summarizes accuracy across distinct spatial query categories, and Figure 6 illustrates performance across varying numbers of objects, clearly highlighting the effectiveness of our interventions.
-Figure 6: Accuracy comparison across different categories of 2DS dataset. We see consistent improvements across categories.
+B.2 How effective can 2DS measure the position embedding sensitivity
+As the preliminary experiments in Section 3 showed that most of the data retain the performance even with random permutation, we can use the same experiments to understand the 2DS results.
+Permutation
+Table 4 show the 2DS with permutation of the input tokens.
+We also report the difference which is used to compute the PSI in Table 2.
+We can see the +Normalized version is more susceptible to the token permutation.
+As the baseline, we also test with the Vicuna-7B to show the model performance is well above the random guess.
+Table 4: For 2DS, we report accuracy on original (Orig.) and permuted (Perm.) token arrangements with difference (Diff.).
+Model 2DS (%)
+Orig. Perm. Diff.
+Vicuna-7B 15.97
+LLaVA 1.5 56.63 33.37 -23.26
++ Norm 59.30 23.00 -36.30
++ Norm + Multi 64.80 29.67 -35.13
+Compression
+Another method to test the 2DS is through compression, which eventually reduce position embeddings effectiveness.
+If the vision encoder can successfully carry over most of the spatial information, then the compression would not hurt performance significantly.
+However, if the position embedding is crucial, then there should be a significant decrease.
+We show the exact same plot but with 2DS through the compression in Figure 9.
+Figure 9: Performance impact of vision token compression on standard benchmarks (GQA, CV-Bench 2D, and POPE).
+Only minor accuracy degradation occurs, even under extreme token compression (down to a single token).
+Here we see a significantly more drop in the 2DS performance. We believe those two measures suggest that position embedding has an important role for the 2DS performance.
+Appendix C Training Details
+We follow the training recipe described in LLaVA 1.5 7B (Liu et al., 2023), consisting of two stages: pretraining and instruction tuning.
+All experiments utilize 8 Nvidia A100 GPUs.
+Learning rates for each training stage are detailed in 5, with all other hyperparameters and datasets identical to (Liu et al., 2023).
+Table 5: Learning rate of training
+Model Pre-train Finetune
+LLaVA 1.5 7B 1e-3 2e-5
++Normalize 1e-3 2e-5
++Normalize+Multilayer 1e-3 2e-5
+Appendix D Analysis of Computational Cost and Inference Time
+In this section, we analyze the computational overhead of our two proposed methods to evaluate their practical efficiency.
+Method 1: +Normalize.
+Our first method, which applies L2 normalization to the final-layer vision tokens, introduces zero additional parameters or floating-point operations (FLOPs) during inference.
+The normalization is a lightweight mathematical operation on the existing embedding vector and its computational cost is negligible.
+Method 2: +Normalize+Multilayer.
+Our second method enriches the visual representation by concatenating features from the intermediate and final layers of the vision encoder.
+This modification affects the initial visual projection layer.
+In the original LLaVA, the projector is an MLP that maps final 1024-dimensional feature vector to 4096-dimensional LLM embedding space.
+While our proposed method concatenates four 1024-dimensional feature vectors and inputs them to the projector, producing 4096-dimensional vector.
+While this moderately increases the number of parameters in this single projection layer, this component represents a very small fraction of the total parameters in the multi-billion parameter VLM.
+Empirical Inference Time.
+To assess the real-world impact on performance, we benchmarked the inference speed of the baseline model against our two proposed methods.
+The experiment was conducted on a single NVIDIA RTX 4090 GPU, measuring the average time to process 500 samples from the GQA dataset validation split.
+All three models are within the same inference time $(0.137s\pm 0.002s)/sample$, empirically there are trivial compute cost differences.
+Appendix E Benchmark Performance
+E.1 2DS Performance
+(a) Accuracy comparison across varying numbers of objects. Our interpretability-informed adjustments yield consistent improvements, especially as spatial complexity increases.
+(b) Accuracy comparison across different categories of 2DS dataset. We see consistent improvements across categories.
+Figure 10: Model performance across increase of objects and different sub-categories
+Table  3 summarizes accuracy across distinct spatial query categories, and Figure  10(b) illustrates performance across varying numbers of objects, clearly highlighting the effectiveness of our interventions.
+Vision Embedding Normalization: Normalization alone improves overall accuracy from 56.63% to 59.30%, strongly supporting our interpretability insight that balancing embedding norms significantly enhances spatial information utilization.
+Notably, improvements are particularly pronounced in color-related spatial relationship queries.
+Normalization + Intermediate-layer Features: This approach achieves the highest overall accuracy (64.80%, +8.17% improvement over baseline), underscoring the importance of explicitly incorporating local spatial details.
+This method substantially enhances accuracy, especially in shape and color combined relative position queries.
+Both methods consistently outperform the original LLaVA 1.5 model, particularly as spatial complexity (the number of objects) increases (Figure  10(a)), validating our hypothesis about the critical roles of embedding normalization and local spatial detail preservation.
+E.2 Results on Standard Benchmarks
+Table 6: Performance (accuracy %) on standard vision-language benchmarks. $\uparrow$ indicates higher is better. Values in parentheses show difference from LLaVA 1.5 baseline.
 Category LLaVA 1.5 + Normalize + Normalize + Multilayer
-Color_abs. $\uparrow$ 79.60 88.00 (+8.40) 87.80 (+8.20)
-Color_rel. $\uparrow$ 37.00 42.40 (+5.40) 47.20 (+10.20)
-Shape_abs. $\uparrow$ 78.60 79.00 (+0.40) 82.20 (+3.60)
-Shape_rel. $\uparrow$ 34.40 32.80 (-1.60) 40.80 (+6.40)
-Shape_color_abs. $\uparrow$ 70.80 71.80 (+1.00) 80.20 (+9.40)
-Shape_color_rel. $\uparrow$ 39.40 41.80 (+2.40) 50.60 (+11.20)
-Overall Acc. $\uparrow$ 56.63 59.30 (+2.67) 64.80 (+8.17)
-Table 2: Spatial reasoning accuracy (%) across 2DS categories. $\uparrow$ indicates higher is better. Values in parentheses show difference from LLaVA 1.5 baseline.
+VQAv2 $\uparrow$ 78.20 78.76 (+0.56) 79.17 (+0.97)
+POPE $\uparrow$ 87.30 87.30 (0.00) 87.70 (+0.40)
+GQA $\uparrow$ 61.46 62.04 (+0.58) 62.52 (+1.05)
+CV-Bench2D $\uparrow$ 56.59 59.91 (+3.32) 58.69 (+2.10)
+Table 7: Performance comparison on CV-Bench 2D benchmarks.
+We report accuracy on original (Orig.) and permuted (Perm.) token arrangements and PSI for each of the three tasks: overall 2D spatial reasoning, counting, and relation.
+We also report evaluation using Vicuna-7B on top without image input but just text input as the random baseline.
+Model CV-Bench 2D (%)
+Overall Count Relation
+Orig. Perm. PSI Orig. Perm. PSI Orig. Perm. PSI
+Vicuna-7B 31.36 12.31 54.46
+LLaVA 1.5 56.59 56.26 0.58 52.07 52.37 -0.58 62.05 60.97 1.74
++ Norm 59.91 53.13 11.32 52.81 51.14 3.16 68.51 55.54 18.93
++ Norm + Multi 58.69 52.99 9.71 52.08 49.49 4.97 66.72 57.23 14.22
+To assess generalization, we evaluate our methods on established benchmarks (Table  6) that are used in recent works (Liu et al., 2023; Tong et al., 2024a; Zhu et al., 2024; Liu et al., 2024).
+Both interpretability-informed interventions yield moderate yet consistent improvements.
+Normalization and Multilayer Normalization consistently achieves small but meaningful gains (up to 3.32%), confirming that improved spatial representation benefits overall reasoning without sacrificing semantic capabilities.
+Notably, normalization shows more pronounced gains in the CV-Bench 2D dataset, particularly for spatial relationship tasks.
+Specifically, we investigate the subcategories of the CV-Bench2D as shown in Table 7.
+We are curious why the permutation drop was not significant.
+As shown in Table 7, we can check the PSI index to see whether the model is sensitive to that task, and Count is not really a spatial task since there are marginal effects when changing spatial information.
+In comparison, we see the Relation task is more spatial, and the +Normalize has a much better performance in that category.
+Appendix F System Prompt Attention
+We visualize the attention portion when inference with and without system prompt for all 3 models in Figure 11.
+We observe that the system prompt absorbs the most attention shares.
+Even when we exclude the system prompt during the calculation, it might skew the probability mass and changes the distribution of a much smaller vision and user input text attentions.
+Therefore we also demonstrate vision and user input text attention inferenced without the system prompt on the right column of the Figure 11.
+Figure 11: Visualization of system prompt, vision, and user input text attention shares.
+Left 2 columns are inferenced with the system prompt, then calculate the portion of the vision and user input text.
+Right column is inferenced without the system prompt.
+We see a much more balanced vision and user input text attention share in the normalized and multilayer version.
+Appendix G Attention Head Visualization
+Figure 12: Visualization of each image’s attention head over layers. We inference without the system prompt. CMB directly calculated from vision and text attention.
+Figure 13: Visualization of each image’s attention head over layers. We inference with the system prompt. CMB take values from vision and text attentions and normalize them with the sum of two.
+Appendix H Attention Visualization
 Figure 14: Visualization of self-attention patterns.
@@
 Each image has entropy on top: lower means more concentrated higher means more diffused
-Vision Embedding Normalization: Normalization alone improves overall accuracy from 56.63% to 59.30%, strongly supporting our interpretability insight that balancing embedding norms significantly enhances spatial information utilization.
-Notably, improvements are particularly pronounced in color-related spatial relationship queries.
-Normalization + Intermediate-layer Features: This approach achieves the highest overall accuracy (64.80%, +8.17% improvement over baseline), underscoring the importance of explicitly incorporating local spatial details.
-This method substantially enhances accuracy, especially in shape and color combined relative position queries.
-Both methods consistently outperform the original LLaVA 1.5 model, particularly as spatial complexity (the number of objects) increases (Figure 5), validating our hypothesis about the critical roles of embedding normalization and local spatial detail preservation.
-6.3 Results on Standard Benchmarks
-Category LLaVA 1.5 + Normalize + Normalize + Multilayer
-VQAv2 $\uparrow$ 78.20 78.76 (+0.56) 79.17 (+0.97)
-POPE $\uparrow$ 87.30 87.30 (0.00) 87.70 (+0.40)
-GQA $\uparrow$ 61.46 62.04 (+0.58) 62.52 (+1.05)
-CV-Bench2D $\uparrow$ 56.59 59.91 (+3.32) 58.69 (+2.10)
-Table 3: Performance (accuracy %) on standard vision-language benchmarks. $\uparrow$ indicates higher is better. Values in parentheses show difference from LLaVA 1.5 baseline.
-To assess generalization, we evaluate our methods on established benchmarks (Table 3) that are used in recent works [23, 32, 41, 24].
-Both interpretability-informed interventions yield moderate yet consistent improvements.
-Normalization and Multilayer Normalization consistently achieves small but meaningful gains (up to 3.32%), confirming that improved spatial representation benefits overall reasoning without sacrificing semantic capabilities.
-Notably, normalization shows more pronounced gains in the CV-Bench 2D dataset, particularly for spatial relationship tasks.
-7 Analysis
-To deepen our interpretability-driven understanding, we analyzed how the combination of embedding normalization and earlier-layer feature extraction influences the attention mechanisms of VLMs.
-7.1 Attention Visualization
 We visualized the self-attention patterns from LLM to interpret the effects of each modification on spatial reasoning.
-Figure 7 provides representative attention maps across four illustrative queries.
+Figure  14 provides representative attention maps across four illustrative queries.
 Each query compares three scenarios: the baseline (Original LLaVA 1.5), normalized embeddings, and normalized embeddings combined with multilayer features.
-We visualized the first text token of the target object in the answer, which are bold and italic words in Figure 7.
-More examples are in Appendix C.
+We visualized the first text token of the target object in the answer, which are bold and italic words in Figure  14.
+More examples are in H.2.
 Baseline (Original LLaVA 1.5): The baseline model shows diffuse and scattered attention patterns, possibly influenced by large-magnitude vision embeddings that overshadow positional cues.
@@
 Consequently, the model selectively focuses attention on regions already identified as informative by intermediate-layer embeddings without the need to check other regions.
-7.2 Interpreting Attention Patterns and Performance
-To quantitatively interpret these observations, we computed the normalized attention entropy (Table 4).
+H.1 Interpreting Attention Patterns and Performance
+To quantitatively interpret these observations, we computed the normalized attention entropy (Table  8).
 We average all text tokens’ attention over the vision tokens to compute the entropy.
 Lower entropy indicates more targeted attention distributions, and high scores mean more diffused attention distributions.
-More examples are in Appendix C.
+More examples are in H.2.
+Table 8: Average attention entropy over different datasets. Lower entropy indicates more selective attention.
 Vision Token Attention Entropy LLaVA 1.5 + Normalize + Normalize + Multilayer
@@
 CV-Bench2D $\downarrow$ 0.80 0.90 0.72
-Table 4: Average attention entropy over different datasets. Lower entropy indicates more selective attention.
 The normalized embeddings exhibit higher entropy compared to the baseline, reflecting a broader and more uniform exploration of spatial regions, yet with a clear focus on the target object.
@@
 In addition, it suggests that the improvements of the Normalize and Multilayer Normalize are two different mechanisms: the former is based on LLM attention, and the later is based on vision encoder spatial information.
-Interpretability Insight No.4.
-This analysis underscores a fundamental interpretability trade-off: richer spatial information embedded earlier in the model pipeline reduces the dependence on explicit positional encoding effect at later stages.
-Consequently, LLM attention becomes sparser yet highly targeted, reflecting enhanced target in spatial information rather than a failure to attend effectively.
 Providing early-layer features from vision encoders potentially offers a shortcut for LLMs to directly leverage spatial cues without extensive positional encoding training.
 Future research should explore strategies to optimally balance intermediate-layer spatial detail and positional encodings.
-8 Discussion
-Our experiments reveal that spatial reasoning in VLMs is often overshadowed by strong semantic cues and disproportionately large vision token norms.
-Normalizing embedding magnitudes and incorporating intermediate-layer vision features partially restore a dorsal-stream-like spatial awareness.
-The most pronounced improvements occur in our synthetic dataset, explicitly designed to emphasize spatial over semantic shortcuts.
-This analysis represents an initial step toward deeper interpretability in VLMs.
-It highlights that not only dataset design but also detailed model analysis is crucial.
-Careful alignment of multi-modal representations, especially when leveraging pretrained models, is essential.
-Our findings provide actionable insights for achieving more effective modality alignment in future research.
-Limitations Despite the promising results, our approach has several limitations: First, the proposed model adjustments are intentionally simple to illustrate our interpretability insights.
-The current normalization method is a brute-force matching to text embedding norms, and it can be overly simplistic.
-Second, there exists a fundamental trade-off regarding where spatial information is encoded.
-Normalization pushes LLMs to utilize positional signals explicitly, while multilayer normalization leverages inherent spatial details from vision encoders Finally, current evaluations focus primarily on 2D spatial reasoning tasks.
-In 3D, the larger magnitude might be more beneficial.
-Future Directions.
-Future work should expand spatial benchmarks to robustly evaluate complex 3D reasoning and explore refined normalization methods that preserve semantic information while still promoting spatial awareness.
-Additionally, architectural modifications explicitly designed for balanced semantic and spatial reasoning could further advance VLM interpretability and effectiveness.
-9 Conclusion
-Through systematic diagnosis and interpretability-driven analysis of VLMs, we propose minimal yet impactful interventions that restore dorsal-stream-like spatial reasoning capabilities.
-Our insights highlight the critical role of embedding norms and layer-specific features, providing a clear interpretability pathway for future model improvements.
-As research increasingly seeks models capable of accurately understanding not only what objects exist but precisely where they are, our contributions lay a foundation for more spatially intelligent, interpretable, and effective VLMs.
-Acknowledgments
-The work is supported by the National Science Foundation (NSF) through Awards #2131186 (CISE-MSI), #1827505 (PFI), and the US Air Force Office of Scientific Research (AFOSR) via Award #FA9550-21-1-0082.
-The work is also supported by a College-wide Research Vision (CRV) Fund from the CCNY Provost’s Office.
-Supplementary Material
-Appendix A Theoretical Analysis of Large-Norm Effects on RoPE
-Problem Setup
-Consider a Transformer attention mechanism [35] employing Rotary Position Embeddings (RoPE)[30].
-Each token $i$ generates a query $\mathbf{q}_{i}\in\mathbb{R}^{d}$ and a key $\mathbf{k}_{i}\in\mathbb{R}^{d}$.
-RoPE applies a position-dependent rotation $R(\cdot)$, transforming these vectors as follows:
-$\mathbf{q}_{i}^{\prime}=R(\mathbf{q}_{i}),\quad\mathbf{k}_{i}^{\prime}=R(\mathbf{k}_{i}).$
-For simplicity, assume $R(\cdot)$ preserves norms (this is standard for RoPE since it only rotates but does not change norm):
-$\|\mathbf{q}_{i}^{\prime}\|=\|\mathbf{q}_{i}\|,\quad\|\mathbf{k}_{i}^{\prime}\|=\|\mathbf{k}_{i}\|.$
-The attention logits are computed using scaled dot-product attention:
-$\mathrm{logit}_{\text{txt},\text{vis}}=\frac{\mathbf{q}_{i}^{\prime}\cdot\mathbf{k}_{j}^{\prime}}{\sqrt{d}}$
-1. Effects of Large Embedding Norms on Attention Weights
-When vision embedding input into LLM directly without model parameter adjustments, we can generally assume the Query and Key of the vision embedding will scale accordingly.
-Consider two sets of embeddings: vision embeddings $\mathbf{q}_{\text{vis}}$, $\mathbf{k}_{\text{vis}}$ and text embeddings $\mathbf{q}_{\text{txt}}$, $\mathbf{k}_{\text{txt}}$.
-Empirically, we observed (as illustrated in the Section  4.1) that vision embeddings typically have substantially larger norms than text embeddings:
-$\|\mathbf{q}_{\text{vis}}\|\approx M\|\mathbf{q}_{\text{txt}}\|,\mathbf{k}_{\text{vis}}\|\approx M\|\mathbf{k}_{\text{txt}}\|,\quad M\gg 1.$
-Empirically, from our study $M$ can reach to the level of $10^{1}$ to $10^{2}$ range.
-Then based on Cauchy-Schwarz inequality we can simply drive following approximations for the attention logits for vision-to-text interactions:
-$\mathrm{logit}_{\text{txt},\text{vis}}=\frac{\mathbf{q}_{\text{txt}}^{\prime}\cdot\mathbf{k}_{\text{vis}}^{\prime}}{\sqrt{d}}\leq\frac{\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{vis}}^{\prime}\|}{\sqrt{d}}\approx\frac{M\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}$
-Similarly, for text-to-text interactions:
-$\mathrm{logit}_{\text{txt},\text{txt}}\approx\frac{\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}$
-Given the scale difference ($M\gg 1$), typically we can see that:
-$\mathrm{logit}_{\text{txt},\text{vis}}\approx\frac{M\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}\gg\frac{\|\mathbf{q}_{\text{txt}}^{\prime}\|\|\mathbf{k}_{\text{txt}}^{\prime}\|}{\sqrt{d}}\approx\mathrm{logit}_{\text{txt},\text{txt}}$
-It would follow that the softmax attention weights from vision tokens towards text tokens become disproportionately large if not adjusted:
-$\alpha_{\text{txt},\text{vis}}$ $=softmax(\frac{\mathbf{q}_{\text{txt}}\cdot\mathbf{k}_{\text{vis}}}{\sqrt{d}})=\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{j}\exp(\mathrm{logit}_{\text{txt},j})}$ (2)
-$=\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{j_{vis}}\exp(\mathrm{logit}_{\text{txt},j_{vis}})+\sum_{j_{txt}}\exp(\mathrm{logit}_{\text{txt},j_{txt}})}$ (3)
-It is easy to see that the terms are dominated by the large $M$ discrepancies between the vision and text embeddings.
-However, practical model training implicitly enforces coherence conditioned primarily on textual tokens, forcing the model to scale down vision-query logits significantly.
-This downscaling results in empirically observed low attention weights for vision tokens.
-[4] observed the attention score of vision tokens are way much smaller compared to the attention weights of text tokens based on the score of the text token, and the score of vision token for most of layers are in $[0,0.2]$ range while text tokens are in $[0.8,1]$ range.
-It is reasonable to believe the model learned to scale down $\mathrm{logit}_{\text{txt},\text{vis}}$ significantly to balance out the large norm.
-2. Impact on RoPE and Positional Encoding
-The attention weight for token $j$ given query token $i$ is defined by the softmax function:
-$\alpha_{\text{txt},\text{vis}}=\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{k}\exp(\mathrm{logit}_{\text{txt},k})}$
-We analyze how sensitively the attention weight $\alpha_{\text{txt},\text{vis}}$ varies with respect to the positional angle $\phi$, introduced by the Rotary Position Embedding (RoPE).
-This sensitivity is given by the derivative:
-$\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\frac{\partial}{\partial\phi}\left(\frac{\exp(\mathrm{logit}_{\text{txt},\text{vis}})}{\sum_{k}\exp(\mathrm{logit}_{\text{txt},k})}\right)$
-Expanding this derivative explicitly with the quotient rule, we can denote $\exp(\mathrm{logit}_{\text{txt},\text{vis}})$ as $u$ and $\sum_{k}\exp(\mathrm{logit}_{\text{txt},k})$ as $v$, then we have:
-$\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\frac{vu^{\prime}-uv^{\prime}}{v^{2}}$
-then simplifying the exponential derivatives and factoring out $\alpha_{\text{txt},\text{vis}}$, we simplify it to:
-$\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\alpha_{\text{txt},\text{vis}}\left(\frac{\partial\mathrm{logit}_{\text{txt},\text{vis}}}{\partial\phi}-\sum_{k}\alpha_{\text{txt},k}\frac{\partial\mathrm{logit}_{\text{txt},k}}{\partial\phi}\right)$
-Key Insight
-From this expression, we observe that the positional sensitivity (derivative) of the attention weight explicitly depends on the attention weight itself ($\alpha_{\text{txt},\text{vis}}$).
-Therefore, when the attention weight $\alpha_{\text{txt},\text{vis}}$ is small (as in the case of vision tokens after significant normalization), the derivative with respect to positional angle $\phi$ is correspondingly reduced, diminishing positional distinctiveness.
-Mathematical Justification: The derivative of attention weights (softmax) w.r.t.
-positional angle $\phi$ is directly proportional to the attention weight itself: $\frac{\partial\alpha_{\text{txt},\text{vis}}}{\partial\phi}=\alpha_{\text{txt},\text{vis}}\left(\frac{\partial\mathrm{logit}_{\text{txt},\text{vis}}}{\partial\phi}-\sum_{k}\alpha_{\text{txt},k}\frac{\partial\mathrm{logit}_{\text{txt},k}}{\partial\phi}\right)$ Thus, lower attention weights inherently reduce positional sensitivity, explaining why downscaled vision embeddings exhibit diminished positional distinctiveness after normalization.
-Appendix B Training Details
-We follow the training recipe described in LLaVA 1.5 7B [23], consisting of two stages: pretraining and instruction tuning.
-All experiments utilize 8 Nvidia A100 GPUs.
-Learning rates for each training stage are detailed in Table 5, with all other hyperparameters and datasets identical to [23].
-Model Pre-train Finetune
-LLaVA 1.5 7B 1e-3 2e-5
-+Normalize 1e-3 2e-5
-+Normalize+Multilayer 1e-3 2e-5
-Table 5: Learning rate of training
-Appendix C Attention Entropy
-We adapt the visualization method from [40] to analyze the LLM attention specifically on vision tokens.
-Attention maps correspond to particular, meaningful text tokens, as labeled above each image in Figure 8.
+H.2 Attention Entropy
+We adapt the visualization method from (Zhang, 2024) to analyze the LLM attention specifically on vision tokens.
+Attention maps correspond to particular, meaningful text tokens, as labeled above each image in 15.
 This is a more refined version compared to average text tokens.